### v2

# Learning to Learn for Few-shot Continual Active Learning

Abstract
Continual learning strives to ensure stability in solving previously seen tasks while demonstrating plasticity in a novel domain. Recent advances in CL are mostly confined to a supervised learning setting, especially in NLP domain. In this work, we consider a few-shot continual active learning (CAL) setting where labeled data are inadequate, and unlabeled data are abundant but with a limited annotation budget. We propose a simple but efficient method, called Meta-Continual Active Learning. Specifically, we employ meta-learning and experience replay to address inter-task confusion and catastrophic forgetting. We further incorporate textual augmentations to ensure generalization. We conduct extensive experiments on benchmark text classification datasets to validate the effectiveness of the proposed method and analyze the effect of different active learning strategies in few-shot CAL setting. Our experimental results demonstrate that random sampling is the best default strategy for active learning and memory sample selection to solve few-shot CAL problems.
Index Terms:
continual learning, meta-learning, active learning, text classification
I Introduction
Continual learning (CL) aims to address the stability-plasticity dilemma in the context of sequential learning. Stability refers to the ability to alleviate the decline in model performance on previously learned tasks, in other words, preventing catastrophic forgetting [1]. Notably, inter-task confusion [2] is one of the major reasons that cause catastrophic forgetting. It is a phenomenon where the model confuses classes from different tasks. Plasticity refers to the ability to adapt to a new task. However, most existing CL methods assume that all data are labeled and sufficient for learning each task, which is impractical.
In real-world scenarios, labeled data are scarce and unlabeled data are abundant. The annotation cost, especially for NLP tasks, is often expensive. To narrow this gap, we consider a more realistic setting for CL, where we provide a small amount of labeled data and a large amount of unlabeled data with limited annotation budget for each CL task. We formulate this problem as a few-shot continual active learning (CAL) process.
Given the success of meta-learning on solving low plasticity, certain studies [3, 4] extend meta-learning to CL setting. In this work, we introduce a simple but efficient method to solve few-shot CAL problems. Specifically, we employ Model-Agnostic Meta-Learning (MAML) framework with memory replay, casting meta-objective as experience replay. In such a way, model learns an optimal initialization that can fast adapt to all seen tasks, thereby enabling fast adaptation while preventing catastrophic forgetting. In addition, we apply consistency regularization via textual augmentations to perform effective learning in the representation space and ensure both intra- and inter-task generalization.
We conduct extensive experiments on CL benchmark datasets from [5], popularized by [6] in lifelong language learning. This collection of datasets includes five text classification datasets from four diverse domains. We validate the effectiveness of our method with different active learning strategies in a 5-shot CAL setup.
The main contributions of this work are threefold:
  • To the best of our knowledge, our paper is the first to address few-shot continual active learning in the NLP domain. The data stream is arbitrary and inherently unpredictable. The label space of a forthcoming task may or may not overlap with that of previous tasks. Hence, we formulate a few-shot continual active learning setting to consider both class-incremental learning (i.e., disjoint label space) and domain-incremental learning (i.e., overlapped label space) in one data sequence.
  • We exploit the advantages of MAML and propose a novel method, namely Meta-Continual Active Learning (Meta-CAL). We deliberately devise meta-objective to align examples from different tasks in current embedding space and subsequently solve problem instances. Especially, we use augmented samples selected from the past as query samples to ensure generalization. As such, the proposed method effectively addresses the problems of inter-task confusion and catastrophic forgetting.
  • In our experiments on various text datasets [5], the proposed method achieves an accuracy of more than 62% while utilizing only 1.6% of seen samples and annotation budgets as low as 500 samples per task. Our study shows that random sampling is the best default strategy for active learning and memory sample selection in the few-shot CAL setting. We observe that random sampling ensures generalization, thereby addressing stability-plasticity dilemma.
The remainder of this paper is organized as follows. We introduce related work in Section II. Preliminaries are described in Section III. Section IV and Section V outline the proposed method. Section VI presents the experimental setup and experimental results. We also include further analysis in Section VII. Finally, Section VIII summarizes this paper and shows the future works.
II Related Work
II-A Learning to learn for CL
In recent years, meta-learning has emerged as an effective learning framework for CL. In particular, the bi-level optimization of meta-learning enables fast adaptation to training data while ensuring generalization across all observed samples. In meta-continuous learning, meta-learning is often combined with memory replay. Meta-MbPA [7] uses MAML to augment episodic memory replay via local adaptation. OML-ER [8] and ANML-ER [8] utilise an online meta-learning model (OML) [9] and a neuromodulated meta-learning (ANML) [10] respectively for effective knowledge transfer through fast adaptation and sparse experience replay. PMR [11] also employs MAML for facilitating episodic memory replay but using a prototypical memory sample selection approach. MER [3] regularizes the objective of experience replay by graident alignment between old and new tasks via a modified Reptile [12]. C-MAML[4] utilizes OML to regulate CL objectives and La-MAML[4] optimizes OML objective through the modulation of per-parameter learning rates. Meta-CL[13] further improves C-MAML and La-MAML by introducing penalty to restrain unnecessary model updates and preserve non-trivial weights for knowledge consolidation. To date, not many models use gradient alignment for continual learning in NLP. In addition, all of these models are employed in a supervised setting. In this work, we focus on aligning meta-learning with CL objectives to provide a solution in a realistic and resource-constrained scenario.
II-B Continual Active Learning
Continual active learning aims to sequentially labels informative data to maximise model performance and solve continual learning problems. It defines a continual learning problem where the available labeled data are insufficient and the annotation budgets are limited. CASA [14] detects new pseudo-domains and selected data from new pseudo-domains for annotation, while it revisits labeled samples to address catastrophic forgetting. [15] proposes a method to address few-shot CAL (FoCAL). They use a Guaussian mixture model (GMM) for active learning and pseudo-rehearsal for CL, bypassing the need to store real past data. However, neither of these methods addresses continual active learning in NLP. CAL-SD [16] tackles NLP tasks, and uses model distillation to augment memory replay with a diversity and uncertainty AL strategies. To date, continual active learning remains understudied, especially in NLP.
III Preliminaries
In this work, we focus on an online task-free class-incremental learning scenario, in which training data stream passes only once without the presence of “task boundaries” [17]. Based on the task-free setting, we formulate the problem of few-shot continual active learning as follows. Assume that training stream consists of $K$ tasks, $\{\mathcal{T}_{1},\mathcal{T}_{2},...,\mathcal{T}_{t},...,\mathcal{T}_{T}\}$. Each task $\mathcal{T}_{t}$ has a small set of labeled instances $\mathcal{D}^{label}_{t}=\{(x_{i},y_{i})\}_{i=1}^{N_{t}\times K}$ and a pool of unlabeled data $\mathcal{D}^{pool}_{t}=\{(u_{i})\}_{i=1}^{|\mathcal{D}^{pool}_{t}|}$. The labeled set consists of $N_{t}$ classes, each with $K$ samples.
III-A Label Space
Based on the label space $\mathcal{Y}$ of tasks, typical CL scenarios have domain-incremental learning, where $\mathcal{Y}_{t}=\bigcup_{i=1}^{T}\mathcal{Y}_{i}$ and class-incremental learning, where $\mathcal{Y}_{t}\cap\mathcal{Y}_{t}^{\prime}=\emptyset,\forall t\neq t^{\prime}$. Since the data stream is arbitrary, the label space $\mathcal{Y}_{k}$ of a task may or may not be disjoint from the other tasks. Hence, we consider both domain-incremental learning and class-incremental learning in one sequence.
III-B Annotation Constraint
We consider each task $\mathcal{T}_{t}$ to have an equal annotation budget $B_{A}$. An acquisition function $a(\cdot)$ dynamically selects informative data points for annotation until acquisition process reaches annotation budget. We denote the selected samples for annotation as $a(\mathcal{D}^{pool}_{t},B_{A})$ and newly annotated sample set as $\mathcal{D}^{new}_{t}$.
III-C Memory Constraint
Replay-based model revisits a small subset of labeled data $\mathcal{D}_{1:1-t}$ to regularize model $f$ while learning $\mathcal{T}_{t}$. We limit the amount of past training data saved in memory $\mathcal{M}$, which should not exceed memory budget threshold $B_{\mathcal{M}}$.
III-D Objectives
Assume we have a learner $f_{\theta}$ and current task $\mathcal{T}_{t}$, our learning objectives are: (a) perform adaptation on $\mathcal{D}_{t}=\mathcal{D}^{label}_{t}\bigcup\mathcal{D}^{new}_{t}$, i.e., plasticity:
$\tilde{\theta}_{t}=\arg\min_{\theta_{t}\in\Theta}\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}[\mathcal{L}\big{(}f_{\theta_{t}}(x),y\big{)}]$ (1)
where $\mathcal{L}_{\mathcal{T}_{t}}$ is the task loss, $\theta_{t}$ is the initial state and $\theta_{t}=\tilde{\theta}_{t-1}$; (b) prevent inter-task confusion and catastrophic forgetting, i.e., stability:
$\min_{\tilde{\theta}_{t}}\frac{1}{|t-1|}\sum_{i=1}^{t-1}\mathbb{E}_{(x,y)\sim\mathcal{D}_{i}}[\mathcal{L}\big{(}f_{\tilde{\theta}_{t}}(x),y\big{)}]$ (2)
III-E Few-shot CAL Challenges
Few-shot CAL problems pose challenges in optimizing neural networks. The main challenges are: (a) quickly adapting the model to the current task $\mathcal{T}_{k}$ while alleviating catastrophic forgetting; (b) addressing inter-task confusion with data scarcity; (c) exploiting efficient active learning strategy that makes the best use of available data to solve continual learning problems.
III-F Active Learning Strategies
In this work, we consider four popular active learning (AL) methods as follows.
Uncertainty
This method samples data $\boldsymbol{x}$ with high uncertainty. The uncertainty is measured by model outputs $\hat{y}$. Least-confidence method [18] evaluates uncertainty by the confidence in predication, where lowest posterior probability indicates greater uncertainty $\alpha_{LC}(\boldsymbol{x},n_{a})=-Pr(\hat{y}|\boldsymbol{x})$. Margin-confidence [19] method considers the confidence margin between two most likely predictions $(\hat{y}_{1},\hat{y}_{2})$, $\alpha_{Marg.}(\boldsymbol{x},n_{a})=-|Pr(\hat{y}_{1}|\boldsymbol{x})-Pr(\hat{y}_{2}|\boldsymbol{x})|$. Small difference indicates high uncertainty. Entropy-based [20] method uses predictive entropy $H$ as the indicator, where higher entropy shows more uncertainty in the posterior probability $\alpha_{Entr.}(\boldsymbol{x},n_{a})=H(Pr(\hat{y}|\boldsymbol{x}))$.
Representative
This method selects data $\boldsymbol{x}$ that are geometrically representative in vector space [21]. In this work, input data $\boldsymbol{x}$ that have the shortest euclidean distance to the centroid of a cluster are representative. KMeans uses the centroid of each cluster, which applies unsupervised learning to partition the data into clusters. The number of clusters is the selection size $n_{a}$. In this work, we also introduce Mean vectors method as a baseline for comparison. It averages representation vectors of each training batch as the centroid.
Diversity
The method chooses data $\boldsymbol{x}$ that are geometrically seen as outliers in vector space [22]. In this work, input data $\boldsymbol{x}$ with the longest Euclidean distance from a centroid are considered diverse. The centroid selections align with those used in the Representative method.
Random
This method randomly samples data $\boldsymbol{x}$ from unlabeled pool $\mathcal{D}^{pool}$.
IV Learning to Learn for CAL
Model-Agnostic Meta-Learning (MAML) [23] is an optimization-based meta-learning approach, often referred to as the “learning-to-learn” algorithm. Learning to learn allows a model to adapt to different data distributions, which can be seen as a form of transfer learning that improves generalization [24]. Therefore, we exploit the MAML framework to facilitate knowledge transfer and generalization across tasks. Moreover, we harness its fast adaption ability to address the challenge of resource scarcity.
IV-A Learning to Fast Adapt
We approximately align the meta-objective with the objectives shown in Eqn. 1 and Eqn. 2 as
$\min_{\theta_{t}}\quad$ $\frac{1}{|t|}\sum_{i=1}^{t}\mathbb{E}_{(x,y)\sim\mathcal{D}_{i}}[\mathcal{L}\big{(}f_{\tilde{\theta}_{t}}(x),y\big{)}]$ (3)
[…]
where $U_{\mathcal{T}_{t},\mathcal{D}_{t}}(\theta_{t})$ is update operation on $\theta_{t}$ using training set $\mathcal{D}_{t}$. Specifically, $U_{\mathcal{T}_{t},\mathcal{D}_{t}}(\theta_{t})$ describes SGD optimization acting on $\theta_{t}$ as
$U_{\mathcal{T}_{t},\mathcal{D}_{t}}(\theta_{t})=\theta_{t}-\alpha\nabla_{\theta_{t}}\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\mathcal{L}(f_{\theta_{t}}(x),y)$ (5)
Hereby, instead of finding finetuned $\tilde{\theta}_{t}$, model $f$ learns an optimal initialization $\theta_{t}$ that effectively adapts to $\mathcal{D}_{1:t}$ with few labeled examples.
IV-B Learning to Continually Learn
In general, the full datasets $\mathcal{D}_{1:t-1}$ are not available while learning $\mathcal{D}_{t}$. Thus, we retrieve past samples from memory set $\mathcal{M}$ for experience replay in meta-objective,
$\min_{\theta_{t}}\quad$ $\sum_{(x,y)\in\mathcal{M}}[\mathcal{L}\big{(}f_{\tilde{\theta}_{t}}(x),y\big{)}]$ (6)
[…]
In such a way, we leverage the selected examples from the past to constrain the learning behaviour of $f$. As a result, it effectively tackles the problems of catastrophic forgetting and inter-task confusion.
IV-C Learning to Generalize
We also exploit data augmentations to enhance model’s ability to generalize. Especially, we apply consistency regularization [25], which is built on the assumption that perturbations of the same input should not affect the output. Inspired by FixMatch [26], we employ two types of data augmentations in meta-training, strong and weak, denoted by $\mathcal{A}(\cdot)$ and $\alpha(\cdot)$, respectively. In contrast to FixMatch, we employ textual augmentations under full supervision to ensure generalization with limited data availability. Specifically, weak and/or strong augmentations are applied in the inner loop to enhance intra-task generalization, while strong augmentations are used in the outer loop to improve both intra- and inter-task generalization. More details are shown in §V-B.
V Model
V-A Model Architecture
Following Online-aware Meta-Learning (OML) [9], the proposed model $f_{\theta}$ consists of a representation learning network $h_{\theta_{\mathrm{e}}}$ with a learnable parameter set $\theta_{\mathrm{e}}$ and a prediction network $g_{\theta_{\mathrm{clf}}}$ with a learnable parameter set $\theta_{\mathrm{clf}}$. The model $f$ is described as $f_{\theta}(x)=g_{\theta_{\mathrm{clf}}}(h_{\theta_{\mathrm{e}}}(x))$. The representation learning network acts as an encoder. The prediction network is a single linear layer followed by a softmax.
V-B Training
Each epoch has $m$ batches of examples instantaneously drawn from data stream. For each task, our model is trained on $\mathcal{D}^{label}$ first, and then on $\mathcal{D}^{pool}$. MAML consists of two optimization loops:
Inner-loop optimization
Inner-loop algorithm performs task-specific learning. We introduce data augmentations as the regularization term to improve intra-task generalization. The inner loop loss for training samples $\mathcal{D}_{i}^{label}$ at time step $i$ is
$\small\begin{split}\mathcal{L}_{\mathbf{inner}}^{D_{i}^{label}}(\theta)&=\sum_{(x,y)\in D_{i}^{label}}[w\mathcal{L}_{CE}\big{(}f_{\theta}(x),y\big{)}\\ &+(1-w)\mathcal{L}_{CE}\big{(}f_{\theta}(\alpha(x)),y\big{)}]\\ \end{split}$ (8)
where $w$ denotes the relative weight and $\mathcal{L}_{CE}$ is the cross-entropy loss. [Annotation process] When the received training batches are unlabeled, we apply acquisition function $a(\mathcal{D}_{i}^{pool},m\cdot n_{a})$ to select informative data points for annotation. The selection size per batch for annotation $n_{a}=\lceil\frac{b\cdot B_{A}}{|\mathcal{D}^{pool}|}\rceil$ where $b$ is the batch size and $B_{A}$ is the annotation budget. Then, the inner-loop loss for newly annotated training batches $\mathcal{D}^{new}_{i}$ is
$\begin{split}\mathcal{L}_{\mathbf{inner}}^{\mathcal{D}^{new}_{i}}(\theta)&=\sum_{(x,y)\in\mathcal{D}^{new}_{i}}[w\mathcal{L}_{CE}\big{(}f_{\theta}(x),y\big{)}\\ &+(1-w)\mathcal{L}_{CE}\big{(}f_{\theta}(\mathcal{A}(x)),f_{\theta}(\alpha(x))\big{)}]\\ \end{split}$ (9)
We use different inner loop loss for already-labeled data $\mathcal{D}_{i}^{label}$ and newly-labeled data $\mathcal{D}^{new}_{i}$. The reason is that newly labeled data may contains more accurate and up-to-date label information. Then, our model performs SGD on parameter set $\theta_{\mathrm{clf}}$ with learning rate $\alpha$ as
$\tilde{\theta}_{\mathrm{clf}}=\theta_{\mathrm{clf}}-\alpha\nabla_{\theta}\mathcal{L}_{\mathbf{inner}}(\theta)$ (10)
[Memory sample selection] We dynamically update memory buffer $\mathcal{M}$ using reservoir sampling to ensure generalization while avoiding overfitting. Reservoir sampling [3] randomly selects a fixed number of training samples without knowing the total number of samples in advance. We use reservoir sampling to select $n_{s}$ examples per class on the incoming data stream $\mathcal{D}^{label}_{i}\bigcup\mathcal{D}^{new}_{i}$ with an equal selection probability for all data seen so far. Note that the current label space $\mathcal{Y}_{i}$ might overlap with previous label spaces, $\mathcal{Y}_{i}\cap(\mathcal{Y}_{1}\cup\mathcal{Y}_{2}\cup\ldots\cup\mathcal{Y}_{i-1})\neq\emptyset$. We automatically update memory samples for $\mathcal{Y}_{i}$ to maintain a fixed amount of memory samples per class.
Input: Initial parameters $\theta=\theta_{\mathrm{e}}\cup\theta_{\mathrm{clf}}$, training set $\mathcal{D}^{train}=\mathcal{D}^{label}\cup\mathcal{D}^{pool}$, memory buffer $\mathcal{M}$, acquisition function $a(\cdot)$, inner-loop learning rate $\alpha$, outer-loop learning rate $\beta$, and memory samples per class $n_{s}$ and annotation budget per task is $B_{A}$.
Output: Trained initial parameters $\theta$ and Memory $\mathcal{M}$
1 for $i=1,2$, … do
2       Receiving $m$ batches of examples, $\mathcal{D}_{i}^{train}$, from the stream
3       if $\mathcal{D}_{i}^{train}$ from a unseen task then
4             Compute selection size per batch for annotation as $n_{a}=\lceil\frac{b\cdot B_{A}}{|\mathcal{D}^{pool}|}\rceil$
5       end if
6      [Inner Loop]
7       // Trained on Labeled data first for each task
8       if $\mathcal{D}_{i}^{train}\subseteq\mathcal{D}^{label}$  then
9             $\mathcal{D}_{i}^{label}\leftarrow\mathcal{D}_{i}^{train}$ and apply weak augmentation on samples from $\mathcal{D}_{i}^{label}$
10             Perform SGD on $\theta_{\mathrm{clf}}$ to minimize Eqn.8
11       else if $\mathcal{D}_{i}^{train}\subseteq\mathcal{D}^{pool}$ then
12             $\mathcal{D}_{i}^{pool}\leftarrow\mathcal{D}_{i}^{train}$ and do annotation per batch as $\mathcal{D}_{i}^{new}\xleftarrow{label}a(\mathcal{D}_{i}^{pool},m\cdot n_{a})$
13             Apply both weak and strong augmentations on samples from $\mathcal{D}_{i}^{new}$
14             Perform SGD on $\theta_{\mathrm{clf}}$ to minimize Eqn.9
15       end if
16      Select memory sample using ReservoirSampling$(\mathcal{D}^{label}_{i}\bigcup\mathcal{D}^{new}_{i},\mathcal{M})$ and update $\mathcal{M}$
17       [Outer Loop]
18       Read ALL examples from $\mathcal{M}$ and apply strong augmentation
19       Perform Adam update on $\theta$ to minimize Eqn.12
20       if all training data are seen then
21             Stop training
22       end if
23
24 end for
Algorithm 1 Continual Active Learning
TABLE I: Average accuracy of four training set orders for sequentially learning five tasks from four domains
Method Save Ratio AL AVG.
RAND REP DIV UNC
MAML-SEQ 100.0% 8.35
$\pm$ 5.1
7.63
$\pm$ 4.8
7.23
$\pm$ 4.1
6.15
$\pm$ 3.9
7.34
$\pm$ 0.9
OML-ER 59.2
$\pm$ 1.9
57.4
$\pm$ 1.7
57.3
$\pm$ 3.1
47.4
$\pm$ 3.3
55.3
$\pm$ 5.4
C-MAML 63.0
$\pm$ 2.9
60.9
$\pm$ 2.8
59.9
$\pm$ 1.0
50.4
$\pm$ 2.1
58.6
$\pm$ 5.6
Meta-CAL (Ours.)
 w/. Prototype 1.6% 57.6
$\pm$ 5.0
57.8
$\pm$ 4.4
57.8
$\pm$ 4.7
42.0
$\pm$ 6.2
53.8
$\pm$ 7.9
 w/. Ring buffer (5 per class) 66.9
$\pm$ 2.4
66.3
$\pm$ 2.2
64.8
$\pm$ 2.5
50.7
$\pm$ 3.6
62.2
$\pm$ 7.7
 w/. Reservoir (default) 66.5
$\pm$ 1.9
65.9
$\pm$ 1.9
64.9
$\pm$ 1.3
53.5
$\pm$ 2.7
62.7
$\pm$ 6.2
FULL 64.1 $\pm$ 1.9
Outer-loop optimization
Outer-loop algorithm optimizes initial parameter set $\theta$ to a setting that $f$ can effectively adapt to $\mathcal{D}_{1:t}$ via a few gradient updates. Due to memory constraints, we only have a few amount of samples from the past. The model reads all examples from memory $\mathcal{M}$. Then, the outer-loop objective is to have $\tilde{\theta}=\theta_{\mathrm{e}}\cup\tilde{\theta}_{\mathrm{clf}}$ generalize well on $\mathcal{D}_{1:t}$ using $\mathcal{M}$, shown in Eqn. 6,
$\begin{split}\mathcal{L}^{\mathcal{M}}_{\mathbf{meta}}(\tilde{\theta})&=\sum_{(x,y)\in\mathcal{M}}[\mathcal{L}_{CE}(f_{\tilde{\theta}}(x),y)]\\ &=\sum_{(x,y)\in\mathcal{M}}[\mathcal{L}_{CE}(g_{\tilde{\theta}_{\mathrm{clf}}}(h_{\theta_{\mathrm{e}}}(x)),y)]\\ \end{split}$ (11)
To improve both intra- and inter-task generalization, we employ strong augmentation to memory samples in meta-objective as
$\mathcal{L}^{\mathcal{M}}_{\mathbf{meta}}(\tilde{\theta})=\sum_{(x,y)\in\mathcal{M}}[\mathcal{L}_{CE}(g_{\tilde{\theta}_{\mathrm{clf}}}(h_{\theta_{\mathrm{e}}}(\mathcal{A}(x)),y)]\$ (12)
To reduce the complexity of the second-order computation in outer loop, we use first-order approximation, namely FOMAML. Then, we use Adam [27] for outer-loop optimization as
$\theta\leftarrow\theta-\beta\nabla_{\tilde{\theta}}\mathcal{L}^{\mathcal{M}}_{\mathbf{meta}}(\tilde{\theta})$ (13)
where $\beta$ is the outer-loop learning rate. Algorithm 1 outlines the complete training procedure.
V-C Testing
The model randomly samples $m$ batches of examples drawn from $\mathcal{M}$ as support set $S$ and performs SGD on these samples to finetune parameter set $\theta_{\mathrm{clf}}$ with learning rate $\alpha$. The inner-loop loss at test is
$\mathcal{L}_{\mathbf{inner}}^{S}(\theta)=\sum_{(x,y)\in S}[\mathcal{L}_{CE}(f_{\theta}(x),y)]\$ (14)
where $S\subseteq\mathcal{M}$. And the optimization process is,
$\tilde{\theta}_{\mathrm{clf}}=\theta_{\mathrm{clf}}-\alpha\nabla_{\theta}\mathcal{L}_{\mathbf{inner}}^{S}(\theta)$ (15)
Then, we output the predication using parameter set $\tilde{\theta}=\theta_{\mathrm{e}}\cup\tilde{\theta}_{\mathrm{clf}}$ on the a test sample $x_{\mathrm{test}}$ as
$\hat{y}_{\mathrm{test}}=f_{\tilde{\theta}}(x_{\mathrm{test}})=g_{\tilde{\theta}_{\mathrm{clf}}}(h_{\theta_{\mathrm{e}}}(x_{\mathrm{test}}))$ (16)
VI Experiments
VI-A Datasets
We use the text classification benchmark datasets from [5], including AGNews (news classification; 4 classes), Yelp (sentiment analysis; 5 classes), Amazon (sentiment analysis; 5 classes), DBpedia (Wikipedia article classification; 14 classes) and Yahoo (questions and answers categorization; 10 classes). This collection contains 5 tasks from 4 different domains, which covers class- and domain-incremental learning in task sequence. We randomly sample 5 labeled instances per class, 10,000 unlabeled instances, and 7,600 test examples from each of the datasets. Following prior studies [7, 8], we concatenate training sets in 4 different orderings as shown in Table II.
TABLE II: Input Datasets Orders
Index Dataset Orders
1 Yelp $\rightarrow$ AGNews $\rightarrow$ DBpedia $\rightarrow$ Amazon $\rightarrow$ Yahoo
2 DBpedia $\rightarrow$ Yahoo $\rightarrow$ AGNews $\rightarrow$ Amazon $\rightarrow$ Yelp
3 Yelp $\rightarrow$ Yahoo $\rightarrow$ Amazon $\rightarrow$ DBpedia $\rightarrow$ AGNews
4 AGNews $\rightarrow$ Yelp $\rightarrow$ Amazon $\rightarrow$ Yahoo $\rightarrow$ DBpedia
VI-B Implementation details
[…]
VI-C Baselines
Baseline model: (a) MAML-SEQ: online FOMAML algorithm. (b) OML-ER [8]: OML with 5% episodic experience replay rate [^3] The default 1% rate shows bad performance. (c) C-MAML [^4] Being modified for NLP tasks [4]: OML with Meta & CL objective alignment + Reservoir Sampling. (d) Meta-CAL (ours.): OML with Meta & CL objective alignment + consistency regularization + reservoir sampling. (e) FULL: supervised C-MAML, trained on full datasets. Memory sample selection: (f) Prototype [11]: select representative samples that are closest to dynamically updated prototypes in representation space. (g) Ring Buffer [30]: use a ‘First-In, First-Out’ scheme to update buffer. (h) Reservoir Sampling [3]: randomly select data with an equal selection probability.
VI-D Evaluation Metrics
Based on the prior work [31], we use three evaluation metrics, i.e., accuracy, backward transfer and forward transfer. Let $R_{k,i}$ be the macro-averaged accuracy evaluated on the test set of the $i$ task after sequentially learning $k$ tasks:
Accuracy
$\mathrm{ACC}_{k}=\frac{1}{k}\sum^{k}_{i=1}R_{k,i}$
Backward transfer
$\mathrm{BWT}_{k}=\frac{1}{k-1}\sum^{k-1}_{i=1}(R_{k,i}-R_{i,i})$
Forward transfer
$\mathrm{FWT}_{k}=\frac{1}{k-1}\sum^{k}_{i=2}(R_{i,i}-R_{0,i}$
TABLE III: Comparison of different AL strategies. We report the average accuracy of training sequence Yelp $\rightarrow$ AGNews $\rightarrow$ DBpedia $\rightarrow$ Amazon $\rightarrow$ Yahoo.
AL Accuracy AVG.
UNC Entropy-based 50.9
$\pm$ 2.7
53.5
Least-confidence 53.5
$\pm$ 2.7
Margin-confidence 56.2
$\pm$ 4.7
DIV Mean-vector 50.7
$\pm$ 1.3
57.8
KMeans 64.9
$\pm$ 1.3
REP Mean-vector 57.9
$\pm$ 2.1
61.9
KMeans 65.9
$\pm$ 1.9
RAND 66.5
$\pm$ 1.9
66.5
(a) Reservoir-RAND(65.8%)
(b) Reservoir-REP (64.7%)
(c) Reservoir-DIV (64.3%)
(d) Reservoir-UNC (53.8%)
(e) RingBuff.-RAND(65.4%)
(f) Proto.-RAND(53.6%)
Figure 1: T-SNE visualization of memory samples in the hidden representation space. Data points with darker colors (purple, violet and pink) represent samples from earlier tasks. The black-circled data points belong to the last task. We provide accuracy for a better comparison.
VI-E Results
In Table I, we compare various models with four AL strategies. These strategies include Random (denoted as RAND), Representative via KMeans (denoted as REP), Diversity via KMeans (denoted as DIV), and Uncertainty via Least-confidence (denoted as UNC). The label budget $B_{A}$ = 2000 examples per task. Each record is the average of three best results from five runs. Save ratio is the ratio of saved samples to the total number of seen samples.
The result shows that our method yields comparable results to the FULL baseline, which is trained on more than 10,000 labeled samples per task. This result indicates our method can effectively select 2,000 informative samples from 10,000 unlabeled data to maximize model performance. Compared to other baselines with the same MAML framework, our method obtains the highest average accuracy among the four commonly-used AL strategies, demonstrating its robustness to different AL approaches. It also indicates Meta-CAL has a great ability of preventing catastrophic forgetting after sequentially learning five tasks.
We also perform paired t-tests. Our model with default setting is significantly better than C-MAML with p-values $<$ 0.04 for all four AL strategies. We also compare AL strategies for the proposed model, Random is significantly better than other strategies with p-values $<$ 0.03.
As for the memory sample selection schemes, Ring Buffer outperforms Reservoir Sampling by less than 1% in RAND and REP. Whereas, Reservoir Sampling exhibits smaller standard deviations, indicating its robustness to training set orders. Prototype Sampling has the worst performance. We perform further analysis in §VII-A.
Both Table I and Table III demonstrate that Random is the best AL strategy for our model. As shown in Table III, for both Diversity- and Representative-based AL strategies, KMeans is the most effective approach. The uncertainty-based methods comparatively perform poorly. CL models intend to revisit past samples that can bring generalization. However, annotating and replaying uncertain samples cannot bring enough generalization, which is not beneficial for knowledge consolidation. Hence, it harms model performance.
VII Further Analysis
We use the training set order [^5] The performance on this order is close to the average performance. Yelp $\rightarrow$ AGNews $\rightarrow$ DBpedia $\rightarrow$ Amazon $\rightarrow$ Yahoo to perform further analysis. In this section, we denote RAND, REP, DIV and UNC as Random, Representative (KMeans), Diversity (KMeans) and Uncertainty (Least-confidence), respectively.
(a) RAND
(b) REP (KMeans)
(c) DIV (KMeans)
(d) UNC (Least-confidence)
Figure 2: Per task accuracy at different learning stages. Darker color indicates higher accuracy. From left to right, the color for each task progressively fades away, indicating forgetting while learning more tasks. UNC shows a lighter color compared to other AL methods, indicating a higher degree of forgetting. The red-outlined box in (a) shows the accuracy on AGNews (Task 2) after learning DBpedia (Task 3).
VII-A Memory Insight
The level of generalization can be inferred from data dispersion. In this section, we investigate the effect of memory samples resulting from active learning and memory sample selection strategies.
AL strategies
As shown in Fig. 1a to 1d, memory data in RAND show a good data dispersion within the last task (i.e., intra-task generalization) and across multiple tasks (i.e., inter-task generalization), resulting in the best accuracy. In contrast, UNC in Fig. 1d shows subpar performance in inter-task generalization. The result validates UNC’s incapability to achieve generalization and preserve knowledge. Both REP and DIV exhibit lower accuracy compared to RAND, with DIV showing an obvious lower sparsity. Therefore, it is important to find an optimal balance between representativeness and diversity to ensure generalization.
Memory sample selection methods
In Fig. 1a, 1e and 1f, the choice of memory sample selection method also affects the model’s performance. Prototype Sampling selects representative memory samples, potentially resulting in relatively less sparsity across multiple tasks. In contrast, Reservoir and Ring Buffer sampling strategies introduce randomness into the memory sample selection process, achieving better inter-task generalization.
Therefore, the randomness introduced by AL and memory sample selection methods can be beneficial in consolidating knowledge by ensuring generalization. Furthermore, it is noteworthy that inter-task confusion does not occur in the learned embedding space. This validates the superiority of our method in addressing confusion between old and new tasks.
VII-B Stability & Plasticity
Next we test the performance of different AL strategies in terms of stability and plasiticity.
(a) BWT (Stability)
(b) FWT (Plasticity)
Figure 3: BWT and FWT for different AL strategies at each learning stage.
Stability
The negative BWT values indicate forgetting. As shown in Fig. 3a, RAND shows the least forgetting, which indicates the best stability on past tasks. However, all methods exhibit a large forgetting after sequentially learning Task 3. Task 3 (DBpedia, 14 classes) has a relatively large label space compared to Task 1 (Yelp, 5 classes) and Task 2 (AGNews, 4 classes). As shown in Fig. 2, the accuracy on DBpedia exceeds the accuracy on other tasks. Since MAML learns high-quality reusable features for fast adaptation [32], it tends to find shared representations that specifically benefit tasks with large label spaces.
Plasticity
A positive FWT demonstrates the model’s ability of zero-shot learning. Task 4 (Amazon) are from the same domain as Task 1 (Yelp). RAND, REP and DIV have a positive FWT value after learning Amazon. The result suggests a positive forward knowledge transfer to the task from the same domain. Moreover, it demonstrates the effectiveness of our model in domain-incremental learning when employed with these three AL strategies.
Overall, RAND shows the best stability while REP shows the best performance in plasticity.
VII-C Effect of Augmentations
We conduct an ablation study to analyze textual augmentations and their key roles in Meta-CAL. We examine the effect of inner-loop augmentation and outer loop augmentation in Table IV. The query samples in the outer-loop constrain the model behaviour. We hypothesize that learning sufficient generalizability from query samples helps retain knowledge and prevent catastrophic forgetting. The results validate this hypothesis, as we observe a significant gain by using outer-loop augmentation to enhance generalization. It improves accuracy by approximately 10% when data availability and label budget are extremely limited. Inner-loop augmentation works well in the extreme case. Nevertheless, the combination of inner- and outer-loop augmentations still improves accuracy.
TABLE IV: Ablation study with different textual augmentation modules of Meta-CAL
AL strategy: 5-shot 1-shot
RAND $B_{A}=2000$ $B_{A}=500$
Meta-CAL (Ours.) 65.7
$\pm$ 0.3
63.5
$\pm$ 0.8
 w/o. inner-loop aug. 65.8
$\pm$ 0.4
62.8
$\pm$ 2.1
 w/o. outer-loop aug. 58.7
$\pm$ 2.3
53.7
$\pm$ 1.0
 w/o. aug. 61.2
$\pm$ 2.1
52.9
$\pm$ 0.9
VII-D Annotation Budgets
As shown in Fig. 4, in contrast to other AL strategies, increasing annotation budgets for UNC degrades model performance. This demonstrates that annotating uncertain samples is not beneficial for knowledge retention. REP achieves the best performance when the label budget is 500, while RAND outperforms other methods in most cases. Consequently, a certain degree of representativeness can aid in knowledge consolidation when annotation budget is extremely limited. In addition, our model achieves more than 62% when the label budget is only 500 samples per task, suggesting its fast adaptation ability.
Figure 4: Performance on different annotation budgets.
VII-E Memory Budgets
We evaluate memory efficiency of our model. Table V compares the performance of three models with the best average accuracy, i.e., our model with Reservoir sampling, our model with Ring Buffer and C-MAML. While all these models use the MAML framework, only the Meta-CAL models employ data augmentation to enhance generalization. The result shows Meta-CAL models outperform C-MAML in all three cases. Notably, Meta-CAL w/. Reservoir can attain more than 50% accuracy by saving only one sample per seen class. As a result, it indicates that data augmentation improves performance.
TABLE V: Accuracy on different memory budgets. The AL strategy is RAND.
Memory Budget Meta-CAL C-MAML
($n_{s}$ per class) Reservoir Ring buffer
1 50.2
$\pm$ 1.9
46.5
$\pm$ 3.4
47.7
$\pm$ 6.0
5 65.7
$\pm$ 0.3
66.4
$\pm$ 0.5
61.2
$\pm$ 2.1
10 67.6
$\pm$ 0.1
68.0
$\pm$ 0.8
63.9
$\pm$ 0.6
Specifically, our model dynamically selects worthwhile unlabeled data to annotate, thereby preventing catastrophic forgetting and maximizing model performance.
VIII Conclusion
This paper considers a realistic scenario, namely few-shot CAL. We devise a novel method, i.e., Meta-Continual Active learning, to address the problem of few-shot CAL. Our model dynamically selects worthwhile unlabeled data to annotate, thereby preventing catastrophic forgetting and maximizing model performance. Specifically, the proposed model employs MAML with experience replay to improve stability and plasticity. We use textual augmentations to enhance intra- and inter-task generalization. In experiments, we examine the proposed model using different active learning strategies. The experimental results demonstrate the effectiveness of our method to solve few-shot continual active learning problems. However, the annotation budget is allocated equally for each task. The annotation budget allocation strategies can be further studied. Furthermore, we can extend our setting to other NLP tasks as future works.

### v3

@@
 We propose a simple but efficient method, called Meta-Continual Active Learning.
+This method sequentially selects the most informative examples from a pool of unlabeled data and requests labels to enhance performance.
 Specifically, we employ meta-learning and experience replay to address inter-task confusion and catastrophic forgetting.
@@
 Index Terms:
-continual learning, meta-learning, active learning, text classification
+continual learning, meta-learning, few-shot learning, active learning, text classification
 I Introduction
@@
 Plasticity refers to the ability to adapt to a new task.
-However, most existing CL methods assume that all data are labeled and sufficient for learning each task, which is impractical.
+However, the majority of existing CL methods presume the availability of labeled data in abundance, deeming it adequate for learning every task.
+Their performance heavily relies on the large quantity and high quality of labeled data, which is impractical.
 In real-world scenarios, labeled data are scarce and unlabeled data are abundant.
-The annotation cost, especially for NLP tasks, is often expensive.
-To narrow this gap, we consider a more realistic setting for CL, where we provide a small amount of labeled data and a large amount of unlabeled data with limited annotation budget for each CL task.
+The cost of annotation, particularly concerning NLP tasks, tends to be prohibitively high.
+To narrow this gap, we contemplate a setting for continual learning that aligns more closely with real-world scenarios.
+Herein, we provide a small amount of labeled data and a large amount of unlabeled data with limited annotation budget for each continual learning task.
 We formulate this problem as a few-shot Continual Active Learning (CAL) process.
+Concretely, model sequentially selects the most worthwhile data from a pool of unlabeled data and requests labels to enhance performance and solve continual learning problems.
 Given the success of meta-learning on solving low plasticity, certain studies [3, 4] extend meta-learning to CL setting.
@@
 The main contributions of this work are threefold:
-  • To the best of our knowledge, our paper is the first to address few-shot continual active learning in the NLP domain.
-The data stream is arbitrary and inherently unpredictable.
-The label space of a forthcoming task may or may not overlap with that of previous tasks.
-Hence, we formulate a few-shot continual active learning setting to consider both class-incremental learning (i.e., disjoint label space) and domain-incremental learning (i.e., overlapped label space) in one data sequence.
+  • Our work tackles few-shot continual active learning challenges in the NLP domain.
+Since the data stream is arbitrary and inherently unpredictable, the label space of a forthcoming task may or may not overlap with that of previous tasks.
+Hence, we formulate a task-free few-shot continual active learning setting to consider both class-incremental learning (i.e., disjoint label space) and domain-incremental learning (i.e., overlapped label space) in one data sequence.
   • We exploit the advantages of MAML and propose a novel method, namely Meta-Continual Active Learning (Meta-CAL).
@@
 Based on the task-free setting, we formulate the problem of few-shot continual active learning as follows.
-Assume that training stream consists of $K$ tasks, $\{\mathcal{T}_{1},\mathcal{T}_{2},...,\mathcal{T}_{t},...,\mathcal{T}_{T}\}$.
+Assume that training stream consists of $T$ tasks, $\{\mathcal{T}_{1},\mathcal{T}_{2},...,\mathcal{T}_{t},...,\mathcal{T}_{T}\}$.
 Each task $\mathcal{T}_{t}$ has a small set of labeled instances $\mathcal{D}^{label}_{t}=\{(x_{i},y_{i})\}_{i=1}^{N_{t}\times K}$ and a pool of unlabeled data $\mathcal{D}^{pool}_{t}=\{(u_{i})\}_{i=1}^{|\mathcal{D}^{pool}_{t}|}$.
@@
 III-A Label Space
-Based on the label space $\mathcal{Y}$ of tasks, typical CL scenarios have domain-incremental learning, where $\mathcal{Y}_{t}=\bigcup_{i=1}^{T}\mathcal{Y}_{i}$ and class-incremental learning, where $\mathcal{Y}_{t}\cap\mathcal{Y}_{t}^{\prime}=\emptyset,\forall t\neq t^{\prime}$.
-Since the data stream is arbitrary, the label space $\mathcal{Y}_{k}$ of a task may or may not be disjoint from the other tasks.
-Hence, we consider both domain-incremental learning and class-incremental learning in one sequence.
+Based on the label space $\mathcal{Y}$ of tasks, typical continual learning scenarios have domain-incremental learning, where $\mathcal{Y}_{t}=\mathcal{Y}_{t}^{\prime},\forall t\neq t^{\prime}$ and class-incremental learning, where $\mathcal{Y}_{t}\cap\mathcal{Y}_{t}^{\prime}=\emptyset,\forall t\neq t^{\prime}$.
+Due to the unforeseen nature of sequential learning, the label space $\mathcal{Y}_{t}$ of a task may or may not be disjoint from the other tasks.
+Hence, we allow $\mathcal{Y}_{t}\cap\mathcal{Y}_{t}^{\prime}\neq\emptyset$ or $\mathcal{Y}_{t}\cap\mathcal{Y}_{t}^{\prime}=\emptyset,\forall t\neq t^{\prime}$ to occur.
 III-B Annotation Constraint
@@
 III-E Few-shot CAL Challenges
-Few-shot CAL problems pose challenges in optimizing neural networks.
-The main challenges are: (a) quickly adapting the model to the current task $\mathcal{T}_{k}$ while alleviating catastrophic forgetting; (b) addressing inter-task confusion with data scarcity; (c) exploiting efficient active learning strategy that makes the best use of available data to solve continual learning problems.
+The few-shot CAL setting poses hurdles for the optimization of neural networks. The main challenges are:
+  • Quickly adapting the model to the current task while alleviating catastrophic forgetting, with limited labeled examples;
+  • Addressing inter-task confusion with data scarcity;
+  • Exploiting an efficient active learning strategy that makes the best use of available data to solve continual learning problems.
 III-F Active Learning Strategies
@@
 V-B Training
-Each epoch has $m$ batches of examples instantaneously drawn from data stream.
+Each episode has $m$ batches of examples instantaneously drawn from data stream.
 For each task, our model is trained on $\mathcal{D}^{label}$ first, and then on $\mathcal{D}^{pool}$.
@@
 VI-B Implementation details
-[…]
+Our example encoder is a pretrained BERTBASE model [28].
+The parameter size of our model is 109M.
+For learning rates, $\alpha=1e^{-3}$ and $\beta=3e^{-5}$.
+The training batch size is 16.
+The number of mini-batches in each episode, $m$ = 5.
+The label budget $B_{A}$ = 2000 examples per task and the memory budget $B_{\mathcal{M}}$ is 5 per class, i.e., $n_{s}=5$.
+For textual augmentation, we swap words randomly as the weak augmentation.
+And, we apply the combination of swapping words randomly, deleting words randomly and substituting words by WordNet’s synonym, as the strong augmentation.
+We use nplug [^1] https://github.com/makcedward/nlpaug 1.1.10, a Python package to implement augmentations.
+All models are executed on Linux platform with 8 Nvidia Tesla A100 GPU and 40 GB of RAM.
+All experiments are performed using PyTorch [^2] https://pytorch.org/ 1.10.0+cu113 [29].
 VI-C Baselines
@@
 (f) Proto.-RAND(53.6%)
-Figure 1: T-SNE visualization of memory samples in the hidden representation space.
+Figure 1: T-SNE visualization of memory samples with different AL strategies.
+A equal dispersion of the data indicates a good memory representation.
+We provide accuracy for a better comparison.
 Data points with darker colors (purple, violet and pink) represent samples from earlier tasks.
 The black-circled data points belong to the last task.
-We provide accuracy for a better comparison.
+The clustering of memory samples from the last task suggests model focuses more on inter-task generalization rather than on intra-task generalization.
 VI-E Results
@@
 In this section, we denote RAND, REP, DIV and UNC as Random, Representative (KMeans), Diversity (KMeans) and Uncertainty (Least-confidence), respectively.
+(a) After Learning Task 1
+(b) After Learning Task 2
+(c) After Learning Task 3
+(d) After Learning Task 4
+(e) After Learning Task 5
+Figure 2: T-SNE visualization of memory samples at different learning stages using training set order Yelp $\rightarrow$ AGNews $\rightarrow$ DBpedia $\rightarrow$ Amazon $\rightarrow$ Yahoo.
+The black-circled data points belong to the last task.
+Data points with darker colors represent samples from earlier tasks, except (d) after learning Task 4.
+Task 4 has the same domain as Task 1.
+Hence, data points with darker colors are belong to the latest task in (d).
 (a) RAND
@@
 Figure 3: Per task accuracy at different learning stages.
-Darker color indicates higher accuracy.
-From left to right, the color for each task progressively fades away, indicating forgetting while learning more tasks.
-UNC shows a lighter color compared to other AL methods, indicating a higher degree of forgetting.
-The red-outlined box in (a) shows the accuracy on AGNews (Task 2) after learning DBpedia (Task 3).
+The dark color indicates high accuracy.
+From left to right, the color for each task progressively fades away, indicating forgetting happens while learning more tasks.
+Note that Yelp and Amazon are from the same domain (sentiment analysis).UNC shows a lighter color compared to other AL methods, indicating a higher degree of forgetting.
+The red-outlined box shows the accuracy on AGNews (Task 2) after learning DBpedia (Task 3).
 VII-A Memory Insight
@@
 This validates the superiority of our method in addressing confusion between old and new tasks.
+Different learning stages
+Fig.
+2 presents the T-SNE visualization of memory samples at different learning stages.
+As show in Figure 2a to 2b, when the proposed model learns a small number of tasks or encounters small label spaces, it focuses on ensuring intra-task generalization.
+However, as the number of seen tasks increases, the model shifts its focus towards ensuring inter-task generalization as shown in 2d to 2e.
+Consequently, we observe a phenomenon wherein the memory samples from the last tasks cluster together.
 VII-B Stability & Plasticity
@@
 As shown in Fig.
-3a, RAND shows the least forgetting, which indicates the best stability on past tasks.
+4a, RAND shows the least forgetting, which indicates the best stability on past tasks.
 However, all methods exhibit a large forgetting after sequentially learning Task 3.
@@
 As shown in Fig.
-2, the accuracy on DBpedia exceeds the accuracy on other tasks.
+3, the accuracy on DBpedia exceeds the accuracy on other tasks.
 Since MAML learns high-quality reusable features for fast adaptation [32], it tends to find shared representations that specifically benefit tasks with large label spaces.
@@
 As shown in Fig.
-4, in contrast to other AL strategies, increasing annotation budgets for UNC degrades model performance.
+5, in contrast to other AL strategies, increasing annotation budgets for UNC degrades model performance.
 This demonstrates that annotating uncertain samples is not beneficial for knowledge retention.
@@
 VIII Conclusion
-This paper considers a realistic scenario, namely few-shot CAL.
-We devise a novel method, i.e., Meta-Continual Active learning, to address the problem of few-shot CAL.
-Our model dynamically selects worthwhile unlabeled data to annotate, thereby preventing catastrophic forgetting and maximizing model performance.
-Specifically, the proposed model employs MAML with experience replay to improve stability and plasticity.
-We use textual augmentations to enhance intra- and inter-task generalization.
-In experiments, we examine the proposed model using different active learning strategies.
-The experimental results demonstrate the effectiveness of our method to solve few-shot continual active learning problems.
-However, the annotation budget is allocated equally for each task.
+This paper considers a realistic continual learning scenario with limited annotation budgets, namely continual active learning.
+We devise a novel method to address this problem.
+Our method dynamically selects worthwhile unlabeled data for annotation to maximize model performance and address continual learning problems.
+Specifically, it applies MAML with memory replay and consistency regularization to enable efficient learning with limited memory and annotation budgets while mitigating catastrophic forgetting.
+In experiments, we analyzed the effect of various active learning strategies and memory sample selection methods on model performance.
+The experimental results shows our method achieves a great performance in a 5-shot continual active learning setup.
+Furthermore, we demonstrate that random sampling is still the best default strategy for both active learning and memory sample selection to solve 5-shot continual active learning problems.
+However, we only evaluate our method in a 5-shot continual active learning scenario.
+In the future work, we can extend our evaluation to a more realistic scenario where the amount of labeled data is {100, 1000, 5000}.
+In addition, the annotation budget is allocated equally for each task.
 The annotation budget allocation strategies can be further studied.
-Furthermore, we can extend our setting to other NLP tasks as future works.

### v4

@@
-Learning to Learn for Few-shot Continual Active Learning
+
 Abstract
 Continual learning strives to ensure stability in solving previously seen tasks while demonstrating plasticity in a novel domain.
-Recent advances in CL are mostly confined to a supervised learning setting, especially in NLP domain.
-In this work, we consider a few-shot continual active learning (CAL) setting where labeled data are inadequate, and unlabeled data are abundant but with a limited annotation budget.
-We propose a simple but efficient method, called Meta-Continual Active Learning.
-This method sequentially selects the most informative examples from a pool of unlabeled data and requests labels to enhance performance.
+Recent advances in continual learning are mostly confined to a supervised learning setting, especially in NLP domain.
+In this work, we consider a few-shot continual active learning setting where labeled data are inadequate, and unlabeled data are abundant but with a limited annotation budget.
+We exploit meta-learning and propose a method, called Meta-Continual Active Learning.
+This method sequentially queries the most informative examples from a pool of unlabeled data for annotation to enhance task-specific performance and tackle continual learning problems through meta-objective.
 Specifically, we employ meta-learning and experience replay to address inter-task confusion and catastrophic forgetting.
-We further incorporate textual augmentations to ensure generalization.
-We conduct extensive experiments on benchmark text classification datasets to validate the effectiveness of the proposed method and analyze the effect of different active learning strategies in few-shot CAL setting.
-Our experimental results demonstrate that random sampling is the best default strategy for active learning and memory sample selection to solve few-shot CAL problems.
-Index Terms:
-continual learning, meta-learning, few-shot learning, active learning, text classification
-I Introduction
+We further incorporate textual augmentations to avoid memory over-fitting caused by experience replay and sample queries, thereby ensuring generalization.
+We conduct extensive experiments on benchmark text classification datasets from diverse domains to validate the feasibility and effectiveness of meta-continual active learning.
+We also analyze the impact of different active learning strategies on various meta continual learning models.
+The experimental results demonstrate that introducing randomness into sample selection is the best default strategy for maintaining generalization in meta-continual learning framework.
+keywords:
+continual learning, meta-learning, active learning, few-shot learning, text classification
+1 Introduction
 Continual Learning (CL) aims to address the stability-plasticity dilemma in the context of sequential learning.
 Stability refers to the ability to alleviate the decline in model performance on previously learned tasks, in other words, preventing catastrophic forgetting [22].
-Notably, inter-task confusion [2] is one of the major reasons that cause catastrophic forgetting.
+Notably, inter-task confusion [15] is one of the major reasons that causes catastrophic forgetting.
 It is a phenomenon where the model confuses classes from different tasks.
@@
 Their performance heavily relies on the large quantity and high quality of labeled data, which is impractical.
-In real-world scenarios, labeled data are scarce and unlabeled data are abundant.
-The cost of annotation, particularly concerning NLP tasks, tends to be prohibitively high.
-To narrow this gap, we contemplate a setting for continual learning that aligns more closely with real-world scenarios.
-Herein, we provide a small amount of labeled data and a large amount of unlabeled data with limited annotation budget for each continual learning task.
-We formulate this problem as a few-shot Continual Active Learning (CAL) process.
-Concretely, model sequentially selects the most worthwhile data from a pool of unlabeled data and requests labels to enhance performance and solve continual learning problems.
+In real-world scenarios, labeled data are scarce while unlabeled data are abundant.
+The cost of annotation, particularly in the field of Natural Language Processing (NLP), tends to be prohibitively high.
+To narrow this gap, we address the problems in a setting that more closely aligns with real-world scenarios, namely few-shot Continual Active Learning (CAL) [2].
+In this setting, only a small subset of labeled data is provided for each task with a limited annotation budget.
+In such a way, model should sequentially select the most worthwhile examples from a pool of unlabeled data and requests labels to enhance performance and simultaneously solve continual learning problems.
+Introducing active learning into the scheme can be challenging as active learning techniques are typically designed to query from a static data distribution.
+They may not be able to dynamically capture most relevant samples for queries to prevent catastrophic forgetting.
+Replay-based methods have been shown to be particularly effective for NLP tasks [37].
+These methods often retain a certain amount of past samples to prevent catastrophic forgetting.
+Consequently, they are prone to memory over-fitting, especially when the availability of labeled examples is limited.
+However, active learning can further escalate the problem by biasing the sample selection towards the memory set.
+Hence, integrating active learning techniques into CL model can be quite challenging.
 Given the success of meta-learning on solving low plasticity, certain studies [29, 14] extend meta-learning to CL setting.
-In this work, we introduce a simple but efficient method to solve few-shot CAL problems.
-Specifically, we employ Model-Agnostic Meta-Learning (MAML) framework with memory replay, casting meta-objective as experience replay.
-In such a way, model learns an optimal initialization that can fast adapt to all seen tasks, thereby enabling fast adaptation while preventing catastrophic forgetting.
-In addition, we apply consistency regularization via textual augmentations to perform effective learning in the representation space and ensure both intra- and inter-task generalization.
-We conduct extensive experiments on CL benchmark datasets from [5], popularized by [6] in lifelong language learning.
+In this work, we exploit the advantages of meta-learning and uses Model-Agnostic Meta-Learning (MAML) [13] framework to solve few-shot CAL problems, called Meta-Continual Active Learning.
+By applying active learning for task-specific tuning and casting meta-objective as experience replay, the learning objective is deliberately formulated to learn an optimal or a suboptimal initial model state that can rapidly adapt to a balanced subset of all encountered task.
+Thereby, this method enables fast adaptation while preventing catastrophic forgetting in few-shot learning.
+In addition, we apply consistency regularization via textual augmentations to address memory overfitting problems that inherent in replay-based method and exacerbated by active learning acquisition.
+We conduct extensive experiments on benchmark datasets from [41], popularized by [12] in lifelong language learning.
 This collection of datasets includes five text classification datasets from four diverse domains.
-We validate the effectiveness of our method with different active learning strategies in a 5-shot CAL setup.
-The main contributions of this work are threefold:
-  • Our work tackles few-shot continual active learning challenges in the NLP domain.
-Since the data stream is arbitrary and inherently unpredictable, the label space of a forthcoming task may or may not overlap with that of previous tasks.
-Hence, we formulate a task-free few-shot continual active learning setting to consider both class-incremental learning (i.e., disjoint label space) and domain-incremental learning (i.e., overlapped label space) in one data sequence.
-  • We exploit the advantages of MAML and propose a novel method, namely Meta-Continual Active Learning (Meta-CAL).
-We deliberately devise meta-objective to align examples from different tasks in current embedding space and subsequently solve problem instances.
-Especially, we use augmented samples selected from the past as query samples to ensure generalization.
-As such, the proposed method effectively addresses the problems of inter-task confusion and catastrophic forgetting.
-  • In our experiments on various text datasets [5], the proposed method achieves an accuracy of more than 62% while utilizing only 1.6% of seen samples and annotation budgets as low as 500 samples per task.
-Our study shows that random sampling is the best default strategy for active learning and memory sample selection in the few-shot CAL setting.
-We observe that random sampling ensures generalization, thereby addressing stability-plasticity dilemma.
-The remainder of this paper is organized as follows.
-We introduce related work in Section II.
-Preliminaries are described in Section III.
-Section IV and Section V outline the proposed method.
-Section VI presents the experimental setup and experimental results.
-We also include further analysis in Section VII.
-Finally, Section VIII summarizes this paper and shows the future works.
-II Related Work
-II-A Learning to learn for CL
+We demonstrate the effectiveness of the proposed framework in a 5-shot CAL setup.
+This paper also examines how various active learning approaches impact the performance of meta continual learning models.
+The main contributions of this paper are fourfold:
+  • Leveraging the strengths of meta-learning, we introduce an optimization-based method, namely Meta-Continual Active Learning (Meta-CAL).
+This method reformulates the meta-objective such that it learns an optimal or a suboptimal initial model state that can effectively adapt to all seen tasks.
+Thereby, it provides a solution to inter-task confusion and catastrophic forgetting even with limited availability of labeled samples.
+  • We integrate active learning into the proposed framework to enhance task-specific tuning.
+This allows the model to dynamically and selectively query the most informative samples from a pool of unlabeled data, thereby improving performance in a resource-constrained scenario.
+  • To address inevitable memory overfitting problems caused by experience replay and active learning, we apply consistency regularization to meta examples via data augmentations.
+This further ensures intra- and inter-task generalization.
+  • In the experiments, the proposed method achieves an accuracy of more than 62% while utilizing only 1.6% of the past samples and maintaining annotation budgets as low as 500 samples for each task.
+The results demonstrate the feasibility and effectiveness of meta-continual active learning.
+Furthermore, we observe that random sampling facilitates generalization in meta continual learning, thereby addressing the stability-plasticity dilemma.
+2 Related Work
+2.1 Continual Learning
+Existing approaches can be categorised into three main mainstreams, i.e., regularization-based methods [20, 21], replay-based methods [12, 17] and architecture-based methods [3, 40].
+Regularization-based methods often impose a penalty or regularization term to the loss function.
+These methods often penalise changes in non-trivial parameters or constrain the variations in gradients learned from previous tasks.
+However, most methods for NLP show a preference for replay-based approaches [37] to avoid unexpected output from tuning the parameters of deep neural networks [38].
+Replay-based methods, also known as rehearsal-based methods or memory-based methods, involves revisiting a small amount of past samples (i.e., experience replay) or generating pseudo past samples while adapting to new domain.
+The popular retrieval schemes for experience replay are random [8, 29, 12, 18], K-Means [38, 16] and Mean-of-Feature [28, 9].
+Architecture-based methods dynamically change the model architecture to learn a new task.
+In general, these methods preserve or partially preserve past fine-tuned parameters and introduce task-specific parameters for the new domain.
+However, it is challenging to manage the scale of the model since the parameter size continuously accumulates with the increase in the number of seen tasks.
+2.2 Meta Continual Learning
 In recent years, meta-learning has emerged as an effective learning framework for CL.
@@
 In this work, we focus on aligning meta-learning with CL objectives to provide a solution in a realistic and resource-constrained scenario.
-II-B Continual Active Learning
+2.3 Continual Active Learning
 Continual active learning aims to sequentially labels informative data to maximise model performance and solve continual learning problems.
@@
 CASA [27] detects new pseudo-domains and selected data from new pseudo-domains for annotation, while it revisits labeled samples to address catastrophic forgetting.
-[15] proposes a method to address few-shot CAL (FoCAL).
+[2] propose a method to address few-shot CAL (FoCAL).
 They use a Guaussian mixture model (GMM) for active learning and pseudo-rehearsal for CL, bypassing the need to store real past data.
@@
 To date, continual active learning remains understudied, especially in NLP.
-III Preliminaries
-In this work, we focus on an online task-free class-incremental learning scenario, in which training data stream passes only once without the presence of “task boundaries” [17].
+3 Preliminaries
+In this work, we focus on task-free class-incremental learning scenario, where training data stream passes only once without the presence of “task boundaries” [34].
 Based on the task-free setting, we formulate the problem of few-shot continual active learning as follows.
 Assume that training stream consists of $T$ tasks, $\{\mathcal{T}_{1},\mathcal{T}_{2},...,\mathcal{T}_{t},...,\mathcal{T}_{T}\}$.
-Each task $\mathcal{T}_{t}$ has a small set of labeled instances $\mathcal{D}^{label}_{t}=\{(x_{i},y_{i})\}_{i=1}^{N_{t}\times K}$ and a pool of unlabeled data $\mathcal{D}^{pool}_{t}=\{(u_{i})\}_{i=1}^{|\mathcal{D}^{pool}_{t}|}$.
-The labeled set consists of $N_{t}$ classes, each with $K$ samples.
-III-A Label Space
+Each task $\mathcal{T}_{t}$ contains $N_{t}$-way $K$-shot set $\mathcal{D}^{label}_{t}=\{(x_{i},y_{i})\}_{i=1}^{N_{t}\times K}$ and a pool of unlabeled data $\mathcal{D}^{pool}_{t}=\{(u_{i})\}_{i=1}^{|\mathcal{D}^{pool}_{t}|}$.
+3.1 Label Space
 Based on the label space $\mathcal{Y}$ of tasks, typical continual learning scenarios have domain-incremental learning, where $\mathcal{Y}_{t}=\mathcal{Y}_{t}^{\prime},\forall t\neq t^{\prime}$ and class-incremental learning, where $\mathcal{Y}_{t}\cap\mathcal{Y}_{t}^{\prime}=\emptyset,\forall t\neq t^{\prime}$.
@@
 Hence, we allow $\mathcal{Y}_{t}\cap\mathcal{Y}_{t}^{\prime}\neq\emptyset$ or $\mathcal{Y}_{t}\cap\mathcal{Y}_{t}^{\prime}=\emptyset,\forall t\neq t^{\prime}$ to occur.
-III-B Annotation Constraint
+3.2 Annotation Constraint
 We consider each task $\mathcal{T}_{t}$ to have an equal annotation budget $B_{A}$.
-An acquisition function $a(\cdot)$ dynamically selects informative data points for annotation until acquisition process reaches annotation budget.
+An acquisition function $a(\cdot)$ dynamically query informative data points for annotation until acquisition process reaches annotation budget.
 We denote the selected samples for annotation as $a(\mathcal{D}^{pool}_{t},B_{A})$ and newly annotated sample set as $\mathcal{D}^{new}_{t}$.
-III-C Memory Constraint
+3.3 Memory Constraint
 Replay-based model revisits a small subset of labeled data $\mathcal{D}_{1:1-t}$ to regularize model $f$ while learning $\mathcal{T}_{t}$.
-We limit the amount of past training data saved in memory $\mathcal{M}$, which should not exceed memory budget threshold $B_{\mathcal{M}}$.
-III-D Objectives
-Assume we have a learner $f_{\theta}$ and current task $\mathcal{T}_{t}$, our learning objectives are: (a) perform adaptation on $\mathcal{D}_{t}=\mathcal{D}^{label}_{t}\bigcup\mathcal{D}^{new}_{t}$, i.e., plasticity:
+We limit the amount of past training data saved in memory buffer $\mathcal{M}$, which should not exceed memory budget threshold $B_{\mathcal{M}}$.
+3.4 Objectives
+Assume we have a learner $f_{\theta}$ and current task $\mathcal{T}_{t}$, the learning objectives are: (a) perform adaptation on $\mathcal{D}_{t}=\mathcal{D}^{label}_{t}\bigcup\mathcal{D}^{new}_{t}$, i.e., plasticity:
 $\tilde{\theta}_{t}=\arg\min_{\theta_{t}\in\Theta}\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}[\mathcal{L}\big{(}f_{\theta_{t}}(x),y\big{)}]$ (1)
-where $\mathcal{L}_{\mathcal{T}_{t}}$ is the task loss, $\theta_{t}$ is the initial state and $\theta_{t}=\tilde{\theta}_{t-1}$; (b) prevent inter-task confusion and catastrophic forgetting, i.e., stability:
+where $\mathcal{L}_{\mathcal{T}_{t}}$ is the task loss, $\theta_{t}$ is the initial state and $\theta_{t}=\tilde{\theta}_{t-1}$; (b) prevent inter-task confusion and catastrophic forgetting of prior tasks, i.e., stability:
 $\min_{\tilde{\theta}_{t}}\frac{1}{|t-1|}\sum_{i=1}^{t-1}\mathbb{E}_{(x,y)\sim\mathcal{D}_{i}}[\mathcal{L}\big{(}f_{\tilde{\theta}_{t}}(x),y\big{)}]$ (2)
-III-E Few-shot CAL Challenges
-The few-shot CAL setting poses hurdles for the optimization of neural networks. The main challenges are:
-  • Quickly adapting the model to the current task while alleviating catastrophic forgetting, with limited labeled examples;
-  • Addressing inter-task confusion with data scarcity;
-  • Exploiting an efficient active learning strategy that makes the best use of available data to solve continual learning problems.
-III-F Active Learning Strategies
+3.5 Active Learning Strategies
 In this work, we consider four popular active learning (AL) methods as follows.
-Uncertainty
+3.5.1 Uncertainty
 This method samples data $\boldsymbol{x}$ with high uncertainty.
@@
 Entropy-based [32] method uses predictive entropy $H$ as the indicator, where higher entropy shows more uncertainty in the posterior probability $\alpha_{Entr.}(\boldsymbol{x},n_{a})=H(Pr(\hat{y}|\boldsymbol{x}))$.
-Representative
+3.5.2 Representative
 This method selects data $\boldsymbol{x}$ that are geometrically representative in vector space [33].
@@
 It averages representation vectors of each training batch as the centroid.
-Diversity
+3.5.3 Diversity
 The method chooses data $\boldsymbol{x}$ that are geometrically seen as outliers in vector space [23].
@@
 The centroid selections align with those used in the Representative method.
-Random
+3.5.4 Random
 This method randomly samples data $\boldsymbol{x}$ from unlabeled pool $\mathcal{D}^{pool}$.
-IV Learning to Learn for CAL
+4 Learning to Learn for CAL
 Model-Agnostic Meta-Learning (MAML) [13] is an optimization-based meta-learning approach, often referred to as the “learning-to-learn” algorithm.
@@
 Moreover, we harness its fast adaption ability to address the challenge of resource scarcity.
-IV-A Learning to Fast Adapt
-We approximately align the meta-objective with the objectives shown in Eqn. 1 and Eqn. 2 as
+4.1 Learning to Fast Adapt
+We approximately align the meta-objective with the objectives shown in Eqn. 1 and Eqn. 2 as follows,
 $\min_{\theta_{t}}\quad$ $\frac{1}{|t|}\sum_{i=1}^{t}\mathbb{E}_{(x,y)\sim\mathcal{D}_{i}}[\mathcal{L}\big{(}f_{\tilde{\theta}_{t}}(x),y\big{)}]$ (3)
-[…]
+s.t. $\tilde{\theta}_{t}=U_{\mathcal{T}_{t},\mathcal{D}_{t}}(\theta_{t})$
 where $U_{\mathcal{T}_{t},\mathcal{D}_{t}}(\theta_{t})$ is update operation on $\theta_{t}$ using training set $\mathcal{D}_{t}$.
-Specifically, $U_{\mathcal{T}_{t},\mathcal{D}_{t}}(\theta_{t})$ describes SGD optimization acting on $\theta_{t}$ as
+Specifically, $U_{\mathcal{T}_{t},\mathcal{D}_{t}}(\theta_{t})$ describes optimization acting on $\theta_{t}$ as
 $U_{\mathcal{T}_{t},\mathcal{D}_{t}}(\theta_{t})=\theta_{t}-\alpha\nabla_{\theta_{t}}\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\mathcal{L}(f_{\theta_{t}}(x),y)$ (4)
-Hereby, instead of finding finetuned $\tilde{\theta}_{t}$, model $f$ learns an optimal initialization $\theta_{t}$ that effectively adapts to $\mathcal{D}_{1:t}$ with few labeled examples.
-IV-B Learning to Continually Learn
+Hereby, instead of finding tuned $\tilde{\theta}_{t}$, model $f$ learns an optimal initialization $\theta_{t}$ that can effectively adapt to $\mathcal{D}_{1:t}$ with few labeled examples.
+4.2 Learning to Continually Learn
 In general, the full datasets $\mathcal{D}_{1:t-1}$ are not available while learning $\mathcal{D}_{t}$.
-Thus, we retrieve past samples from memory set $\mathcal{M}$ for experience replay in meta-objective,
+Thus, we retrieve a small subset of past samples $\mathcal{M}$ for experience replay in meta-objective,
 $\min_{\theta_{t}}\quad$ $\sum_{(x,y)\in\mathcal{M}}[\mathcal{L}\big{(}f_{\tilde{\theta}_{t}}(x),y\big{)}]$ (5)
-[…]
+s.t. $\tilde{\theta}_{t}=U_{\mathcal{T}_{t},\mathcal{D}_{t}}(\theta_{t})$
+$=\theta_{t}-\alpha\nabla_{\theta_{t}}\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\mathcal{L}(f_{\theta_{t}}(x),y)$
 In such a way, we leverage the selected examples from the past to constrain the learning behaviour of $f$.
-As a result, it effectively tackles the problems of catastrophic forgetting and inter-task confusion.
-IV-C Learning to Generalize
+As a result, it effectively tackles the problems of inter-task confusion and catastrophic forgetting.
+4.3 Learning to Generalize
 We also exploit data augmentations to enhance model’s ability to generalize.
@@
 Specifically, weak and/or strong augmentations are applied in the inner loop to enhance intra-task generalization, while strong augmentations are used in the outer loop to improve both intra- and inter-task generalization.
-More details are shown in §V-B.
-V Model
-V-A Model Architecture
+More details are shown in §5.2.
+5 Model
+5.1 Model Architecture
 Following Online-aware Meta-Learning (OML) [19], the proposed model $f_{\theta}$ consists of a representation learning network $h_{\theta_{\mathrm{e}}}$ with a learnable parameter set $\theta_{\mathrm{e}}$ and a prediction network $g_{\theta_{\mathrm{clf}}}$ with a learnable parameter set $\theta_{\mathrm{clf}}$.
@@
 The prediction network is a single linear layer followed by a softmax.
-V-B Training
+5.2 Training
 Each episode has $m$ batches of examples instantaneously drawn from data stream.
@@
 MAML consists of two optimization loops:
-Inner-loop optimization
-Inner-loop algorithm performs task-specific learning.
+5.2.1 Inner-Loop Optimization
+Inner-loop algorithm performs task-specific tuning.
 We introduce data augmentations as the regularization term to improve intra-task generalization.
 The inner loop loss for training samples $\mathcal{D}_{i}^{label}$ at time step $i$ is
-$\small\begin{split}\mathcal{L}_{\mathbf{inner}}^{D_{i}^{label}}(\theta)&=\sum_{(x,y)\in D_{i}^{label}}[w\mathcal{L}_{CE}\big{(}f_{\theta}(x),y\big{)}\\ &+(1-w)\mathcal{L}_{CE}\big{(}f_{\theta}(\alpha(x)),y\big{)}]\\ \end{split}$ (8)
+$\small\mathcal{L}_{\mathbf{inner}}^{D_{i}^{label}}(\theta)=\sum_{(x,y)\in D_{i}^{label}}[w\mathcal{L}_{CE}(f_{\theta}(x),y)+(1-w)\mathcal{L}_{CE}\big{(}f_{\theta}(\alpha(x)),y\big{)}]$ (6)
 where $w$ denotes the relative weight and $\mathcal{L}_{CE}$ is the cross-entropy loss.
-[Annotation process] When the received training batches are unlabeled, we apply acquisition function $a(\mathcal{D}_{i}^{pool},m\cdot n_{a})$ to select informative data points for annotation.
+Annotation Process
+When the received training batches are unlabeled, we apply acquisition function $a(\mathcal{D}_{i}^{pool},m\cdot n_{a})$ to select informative data points for annotation.
 The selection size per batch for annotation $n_{a}=\lceil\frac{b\cdot B_{A}}{|\mathcal{D}^{pool}|}\rceil$ where $b$ is the batch size and $B_{A}$ is the annotation budget.
 Then, the inner-loop loss for newly annotated training batches $\mathcal{D}^{new}_{i}$ is
-$\begin{split}\mathcal{L}_{\mathbf{inner}}^{\mathcal{D}^{new}_{i}}(\theta)&=\sum_{(x,y)\in\mathcal{D}^{new}_{i}}[w\mathcal{L}_{CE}\big{(}f_{\theta}(x),y\big{)}\\ &+(1-w)\mathcal{L}_{CE}\big{(}f_{\theta}(\mathcal{A}(x)),f_{\theta}(\alpha(x))\big{)}]\\ \end{split}$ (9)
+$\small\mathcal{L}_{\mathbf{inner}}^{\mathcal{D}^{new}_{i}}(\theta)=\sum_{(x,y)\in\mathcal{D}^{new}_{i}}[w\mathcal{L}_{CE}\big{(}f_{\theta}(x),y\big{)}+(1-w)\mathcal{L}_{CE}\big{(}f_{\theta}(\mathcal{A}(x)),f_{\theta}(\alpha(x))\big{)}]$ (7)
 We use different inner loop loss for already-labeled data $\mathcal{D}_{i}^{label}$ and newly-labeled data $\mathcal{D}^{new}_{i}$.
@@
 $\tilde{\theta}_{\mathrm{clf}}=\theta_{\mathrm{clf}}-\alpha\nabla_{\theta}\mathcal{L}_{\mathbf{inner}}(\theta)$ (8)
-[Memory sample selection] We dynamically update memory buffer $\mathcal{M}$ using reservoir sampling to ensure generalization while avoiding overfitting.
+Memory Sample Selection
+We dynamically update memory buffer $\mathcal{M}$ using reservoir sampling to ensure generalization while avoiding overfitting.
 Reservoir sampling [29] randomly selects a fixed number of training samples without knowing the total number of samples in advance.
 We use reservoir sampling to select $n_{s}$ examples per class on the incoming data stream $\mathcal{D}^{label}_{i}\bigcup\mathcal{D}^{new}_{i}$ with an equal selection probability for all data seen so far.
-Note that the current label space $\mathcal{Y}_{i}$ might overlap with previous label spaces, $\mathcal{Y}_{i}\cap(\mathcal{Y}_{1}\cup\mathcal{Y}_{2}\cup\ldots\cup\mathcal{Y}_{i-1})\neq\emptyset$.
+Note that the current label space $\mathcal{Y}_{i}$ might overlap with previous label spaces.
 We automatically update memory samples for $\mathcal{Y}_{i}$ to maintain a fixed amount of memory samples per class.
-Input: Initial parameters $\theta=\theta_{\mathrm{e}}\cup\theta_{\mathrm{clf}}$, training set $\mathcal{D}^{train}=\mathcal{D}^{label}\cup\mathcal{D}^{pool}$, memory buffer $\mathcal{M}$, acquisition function $a(\cdot)$, inner-loop learning rate $\alpha$, outer-loop learning rate $\beta$, and memory samples per class $n_{s}$ and annotation budget per task is $B_{A}$.
-Output: Trained initial parameters $\theta$ and Memory $\mathcal{M}$
-1 for $i=1,2$, … do
-2       Receiving $m$ batches of examples, $\mathcal{D}_{i}^{train}$, from the stream
-3       if $\mathcal{D}_{i}^{train}$ from a unseen task then
-4             Compute selection size per batch for annotation as $n_{a}=\lceil\frac{b\cdot B_{A}}{|\mathcal{D}^{pool}|}\rceil$
-5       end if
-6      [Inner Loop]
-7       // Trained on Labeled data first for each task
-8       if $\mathcal{D}_{i}^{train}\subseteq\mathcal{D}^{label}$  then
-9             $\mathcal{D}_{i}^{label}\leftarrow\mathcal{D}_{i}^{train}$ and apply weak augmentation on samples from $\mathcal{D}_{i}^{label}$
-10             Perform SGD on $\theta_{\mathrm{clf}}$ to minimize Eqn.8
-11       else if $\mathcal{D}_{i}^{train}\subseteq\mathcal{D}^{pool}$ then
-12             $\mathcal{D}_{i}^{pool}\leftarrow\mathcal{D}_{i}^{train}$ and do annotation per batch as $\mathcal{D}_{i}^{new}\xleftarrow{label}a(\mathcal{D}_{i}^{pool},m\cdot n_{a})$
-13             Apply both weak and strong augmentations on samples from $\mathcal{D}_{i}^{new}$
-14             Perform SGD on $\theta_{\mathrm{clf}}$ to minimize Eqn.9
-15       end if
-16      Select memory sample using ReservoirSampling$(\mathcal{D}^{label}_{i}\bigcup\mathcal{D}^{new}_{i},\mathcal{M})$ and update $\mathcal{M}$
-17       [Outer Loop]
-18       Read ALL examples from $\mathcal{M}$ and apply strong augmentation
-19       Perform Adam update on $\theta$ to minimize Eqn.12
-20       if all training data are seen then
-21             Stop training
-22       end if
-23
-24 end for
 Algorithm 1 Continual Active Learning
-TABLE I: Average accuracy of four training set orders for sequentially learning five tasks from four domains
-Method Save Ratio AL AVG.
+1:Training set $\mathcal{D}^{train}=\mathcal{D}^{label}\cup\mathcal{D}^{pool}$, acquisition function $a(\cdot)$, inner-loop learning rate $\alpha$, outer-loop learning rate $\beta$, memory samples per class $n_{s}$, annotation budget per task $B_{A}$
+2:Trained model with parameters $\theta=\theta_{\mathrm{e}}\cup\theta_{\mathrm{clf}}$ and memory buffer $\mathcal{M}$
+3:for $i=1,2,\ldots$ do
+4:    Receiving $m$ batches of examples, $\mathcal{D}_{i}^{train}$, from the stream
+5:    if $\mathcal{D}_{i}^{train}$ is from an unseen task then
+6:         Compute selection size per batch for annotation as $n_{a}=\lceil\frac{b\cdot B_{A}}{|\mathcal{D}^{pool}|}\rceil$
+7:    end if
+8:    [Inner Loop] $\triangleright$ Trained on Labeled data first for each task
+9:    if $\mathcal{D}_{i}^{train}\subseteq\mathcal{D}^{label}$ then
+10:         $\mathcal{D}_{i}^{label}\leftarrow\mathcal{D}_{i}^{train}$ and apply weak augmentation on samples from $\mathcal{D}_{i}^{label}$
+11:         Perform SGD on $\theta_{\mathrm{clf}}$ to minimize Eqn. 6
+12:    else if $\mathcal{D}_{i}^{train}\subseteq\mathcal{D}^{pool}$ then
+[…]
+14:         Apply both weak and strong augmentations on samples from $\mathcal{D}_{i}^{new}$
+15:         Perform SGD on $\theta_{\mathrm{clf}}$ to minimize Eqn. 7
+16:    end if
+17:    Select memory sample using ReservoirSampling$(\mathcal{D}^{label}_{i}\bigcup\mathcal{D}^{new}_{i},\mathcal{M})$ and update $\mathcal{M}$
+18:    [Outer Loop]
+19:    Read ALL examples from $\mathcal{M}$ and apply strong augmentation
+20:    Perform Adam update on $\theta$ to minimize Eqn. 10
+21:    if all training data are seen then
+22:         Stop training
+23:    end if
+24:end for
+5.2.2 Outer-Loop Optimization
+Outer-loop algorithm optimizes initial parameter set $\theta$ to a setting that $f$ can effectively adapt to $\mathcal{D}_{1:t}$ via a few gradient updates.
+Due to memory constraints, we only have a few amount of samples from the past.
+The model reads all examples from memory buffer $\mathcal{M}$.
+Then, the outer-loop objective is to have $\tilde{\theta}=\theta_{\mathrm{e}}\cup\tilde{\theta}_{\mathrm{clf}}$ generalize well on $\mathcal{D}_{1:t}$ using $\mathcal{M}$, shown in Eqn.
+5,
+$\begin{split}\mathcal{L}^{\mathcal{M}}_{\mathbf{meta}}(\tilde{\theta})&=\sum_{(x,y)\in\mathcal{M}}[\mathcal{L}_{CE}(f_{\tilde{\theta}}(x),y)]\\ &=\sum_{(x,y)\in\mathcal{M}}[\mathcal{L}_{CE}(g_{\tilde{\theta}_{\mathrm{clf}}}(h_{\theta_{\mathrm{e}}}(x)),y)]\\ \end{split}$ (9)
+To improve both intra- and inter-task generalization, we employ strong augmentation to memory samples in meta-objective as
+$\mathcal{L}^{\mathcal{M}}_{\mathbf{meta}}(\tilde{\theta})=\sum_{(x,y)\in\mathcal{M}}[\mathcal{L}_{CE}(g_{\tilde{\theta}_{\mathrm{clf}}}(h_{\theta_{\mathrm{e}}}(\mathcal{A}(x)),y)]\$ (10)
+To reduce the complexity of the second-order computation in outer loop, we use first-order approximation, namely FOMAML. The outer-loop optimization process is
+$\theta\leftarrow\theta-\beta\nabla_{\tilde{\theta}}\mathcal{L}^{\mathcal{M}}_{\mathbf{meta}}(\tilde{\theta})$ (11)
+where $\beta$ is the outer-loop learning rate. Algorithm 1 outlines the complete training procedure.
+5.3 Testing
+The model randomly samples $m$ batches of examples drawn from $\mathcal{M}$ as support set $S$ and performs SGD on these samples to finetune parameter set $\theta_{\mathrm{clf}}$ with learning rate $\alpha$.
+The inner-loop loss at test is
+$\mathcal{L}_{\mathbf{inner}}^{S}(\theta)=\sum_{(x,y)\in S}[\mathcal{L}_{CE}(f_{\theta}(x),y)]\$ (12)
+where $S\subseteq\mathcal{M}$. And the optimization process is,
+$\tilde{\theta}_{\mathrm{clf}}=\theta_{\mathrm{clf}}-\alpha\nabla_{\theta}\mathcal{L}_{\mathbf{inner}}^{S}(\theta)$ (13)
+Then, we output the predication using parameter set $\tilde{\theta}=\theta_{\mathrm{e}}\cup\tilde{\theta}_{\mathrm{clf}}$ on the a test sample $x_{\mathrm{test}}$ as
+$\hat{y}_{\mathrm{test}}=f_{\tilde{\theta}}(x_{\mathrm{test}})=g_{\tilde{\theta}_{\mathrm{clf}}}(h_{\theta_{\mathrm{e}}}(x_{\mathrm{test}}))$ (14)
+6 Experiments
+6.1 Datasets
+We use the text classification benchmark datasets from [41], including AGNews (news classification; 4 classes), Yelp (sentiment analysis; 5 classes), Amazon (sentiment analysis; 5 classes), DBpedia (Wikipedia article classification; 14 classes) and Yahoo (questions and answers categorization; 10 classes).
+This collection contains 5 tasks from 4 different domains, which covers class- and domain-incremental learning in task sequence.
+We randomly sample 5 labeled instances per class, 10,000 unlabeled instances, and 7,600 test examples from each of the datasets.
+Following prior studies [36, 18, 17], we concatenate training sets in 4 different orderings as shown in Table 1.
+Table 1: Input Dataset Orders
+Index Dataset Orders
+1 Yelp $\rightarrow$ AGNews $\rightarrow$ DBpedia $\rightarrow$ Amazon $\rightarrow$ Yahoo
+2 DBpedia $\rightarrow$ Yahoo $\rightarrow$ AGNews $\rightarrow$ Amazon $\rightarrow$ Yelp
+3 Yelp $\rightarrow$ Yahoo $\rightarrow$ Amazon $\rightarrow$ DBpedia $\rightarrow$ AGNews
+4 AGNews $\rightarrow$ Yelp $\rightarrow$ Amazon $\rightarrow$ Yahoo $\rightarrow$ DBpedia
+[…]
+6.2 Implementation details
+Our example encoder is a pretrained BERTBASE model [11].
+The parameter size of our model is 109M.
+For learning rates, $\alpha=1e^{-3}$ and $\beta=3e^{-5}$.
+The training batch size is 16.
+The number of mini-batches in each episode, $m$ = 5.
+The label budget $B_{A}$ = 2000 examples per task and the memory budget $B_{\mathcal{M}}$ is 5 per class, i.e., $n_{s}=5$.
+For textual augmentation, we swap words randomly as the weak augmentation.
+And, we apply the combination of swapping words randomly, deleting words randomly and substituting words by WordNet’s synonym, as the strong augmentation.
+We use nplug [^1] https://github.com/makcedward/nlpaug 1.1.10, a Python package to implement augmentations.
+All models are executed on Linux platform with 8 Nvidia Tesla A100 GPU and 40 GB of RAM.
+All experiments are performed using PyTorch [^2] https://pytorch.org/ 1.10.0+cu113 [26].
+6.3 Baselines
+Baseline model:
+  1. 1.
+    MAML-SEQ: Online FOMAML algorithm.
+  2. 2.
+    OML-ER [18]: OML with 5% episodic experience replay rate [^3] The default 1% rate shows bad performance + reservoir sampling.
+  3. 3.
+    C-MAML [^4] Being modified for NLP tasks [14]: OML with Meta & CL objective alignment + reservoir sampling.
+  4. 4.
+    Meta-CAL (ours): OML with Meta & CL objective alignment + consistency regularization + reservoir sampling.
+  5. 5.
+    FULL: Supervised C-MAML, trained on full datasets.
+Memory sample selection:
+  1. 1.
+    Prototype [17]: Select representative samples that are closest to dynamically updated prototypes in representation space.
+  2. 2.
+    Ring Buffer [7]: Use a ’First-In, First-Out’ scheme to update buffer.
+  3. 3.
+    Reservoir Sampling [29]: Randomly select data with an equal selection probability.
+6.4 Evaluation Metrics
+Based on the prior work [39], we use three comprehensive and widely-used metrics in CL, i.e., accuracy, backward transfer and forward transfer.
+Let $R_{k,i}$ be the macro-averaged accuracy evaluated on the test set of the $i$ task after sequentially learning $t$ tasks,
+Accuracy:
+$\mathrm{ACC}_{t}=\frac{1}{t}\sum^{t}_{i=1}R_{t,i}$
+Overall accuracy is the weighted average accuracy of all seen tasks $\mathcal{T}_{1:T}$.
+Backward transfer (stability evaluation):
+$\mathrm{BWT}_{k}=\frac{1}{k-1}\sum^{k-1}_{i=1}(R_{k,i}-R_{i,i})$
+BWT measures how the updated parameters affect model performance on all previously seen tasks.
+Forward transfer (plasticity evaluation):
+$\mathrm{FWT}_{k}=\frac{1}{k-1}\sum^{k}_{i=2}(R_{i,i}-R_{0,i})$
+FWT quantifies the average impact of all preceding tasks on the current $k$ task.
+6.5 Main Results
+In Table 2, we compare various models with four AL strategies.
+These strategies include Random (denoted as RAND), Representative via KMeans (denoted as REP), Diversity via KMeans (denoted as DIV), and Uncertainty via Least-confidence (denoted as UNC).
+The label budget $B_{A}$ = 2000 examples per task.
+Each record is the average of three best results from five runs.
+Table 2: Average Accuracy of Four Training Set Orders
+Method Save Ratio [^1] The ratio of saved samples to the total number of seen samples AL AVG.
 RAND REP DIV UNC
@@
 $\pm$ 5.6
-Meta-CAL (Ours.)
+Meta-CAL (Ours)
    w/. Prototype 1.6% 57.6
@@
 $\pm$ 7.9
-   w/. Ring buffer (5 per class) 66.9
+   w/. Ring buffer (5 per class) $\mathbf{66.9}$
 $\pm$ 2.4
@@
 $\pm$ 7.7
-   w/. Reservoir (default) 66.5
+   w/. Reservoir [^2] Default memory selection scheme 66.5
 $\pm$ 1.9
@@
 $\pm$ 6.2
-FULL 64.1 $\pm$ 1.9
-Outer-loop optimization
-Outer-loop algorithm optimizes initial parameter set $\theta$ to a setting that $f$ can effectively adapt to $\mathcal{D}_{1:t}$ via a few gradient updates.
-Due to memory constraints, we only have a few amount of samples from the past.
-The model reads all examples from memory $\mathcal{M}$.
-Then, the outer-loop objective is to have $\tilde{\theta}=\theta_{\mathrm{e}}\cup\tilde{\theta}_{\mathrm{clf}}$ generalize well on $\mathcal{D}_{1:t}$ using $\mathcal{M}$, shown in Eqn.
-6,
-$\begin{split}\mathcal{L}^{\mathcal{M}}_{\mathbf{meta}}(\tilde{\theta})&=\sum_{(x,y)\in\mathcal{M}}[\mathcal{L}_{CE}(f_{\tilde{\theta}}(x),y)]\\ &=\sum_{(x,y)\in\mathcal{M}}[\mathcal{L}_{CE}(g_{\tilde{\theta}_{\mathrm{clf}}}(h_{\theta_{\mathrm{e}}}(x)),y)]\\ \end{split}$ (11)
-To improve both intra- and inter-task generalization, we employ strong augmentation to memory samples in meta-objective as
-$\mathcal{L}^{\mathcal{M}}_{\mathbf{meta}}(\tilde{\theta})=\sum_{(x,y)\in\mathcal{M}}[\mathcal{L}_{CE}(g_{\tilde{\theta}_{\mathrm{clf}}}(h_{\theta_{\mathrm{e}}}(\mathcal{A}(x)),y)]\$ (12)
-To reduce the complexity of the second-order computation in outer loop, we use first-order approximation, namely FOMAML. Then, we use Adam [27] for outer-loop optimization as
-$\theta\leftarrow\theta-\beta\nabla_{\tilde{\theta}}\mathcal{L}^{\mathcal{M}}_{\mathbf{meta}}(\tilde{\theta})$ (13)
-where $\beta$ is the outer-loop learning rate. Algorithm 1 outlines the complete training procedure.
-V-C Testing
-The model randomly samples $m$ batches of examples drawn from $\mathcal{M}$ as support set $S$ and performs SGD on these samples to finetune parameter set $\theta_{\mathrm{clf}}$ with learning rate $\alpha$.
-The inner-loop loss at test is
-$\mathcal{L}_{\mathbf{inner}}^{S}(\theta)=\sum_{(x,y)\in S}[\mathcal{L}_{CE}(f_{\theta}(x),y)]\$ (14)
-where $S\subseteq\mathcal{M}$. And the optimization process is,
-$\tilde{\theta}_{\mathrm{clf}}=\theta_{\mathrm{clf}}-\alpha\nabla_{\theta}\mathcal{L}_{\mathbf{inner}}^{S}(\theta)$ (15)
-Then, we output the predication using parameter set $\tilde{\theta}=\theta_{\mathrm{e}}\cup\tilde{\theta}_{\mathrm{clf}}$ on the a test sample $x_{\mathrm{test}}$ as
-$\hat{y}_{\mathrm{test}}=f_{\tilde{\theta}}(x_{\mathrm{test}})=g_{\tilde{\theta}_{\mathrm{clf}}}(h_{\theta_{\mathrm{e}}}(x_{\mathrm{test}}))$ (16)
-VI Experiments
-VI-A Datasets
-We use the text classification benchmark datasets from [5], including AGNews (news classification; 4 classes), Yelp (sentiment analysis; 5 classes), Amazon (sentiment analysis; 5 classes), DBpedia (Wikipedia article classification; 14 classes) and Yahoo (questions and answers categorization; 10 classes).
-This collection contains 5 tasks from 4 different domains, which covers class- and domain-incremental learning in task sequence.
-We randomly sample 5 labeled instances per class, 10,000 unlabeled instances, and 7,600 test examples from each of the datasets.
-Following prior studies [7, 8], we concatenate training sets in 4 different orderings as shown in Table II.
-TABLE II: Input Datasets Orders
-Index Dataset Orders
-1 Yelp $\rightarrow$ AGNews $\rightarrow$ DBpedia $\rightarrow$ Amazon $\rightarrow$ Yahoo
-2 DBpedia $\rightarrow$ Yahoo $\rightarrow$ AGNews $\rightarrow$ Amazon $\rightarrow$ Yelp
-3 Yelp $\rightarrow$ Yahoo $\rightarrow$ Amazon $\rightarrow$ DBpedia $\rightarrow$ AGNews
-4 AGNews $\rightarrow$ Yelp $\rightarrow$ Amazon $\rightarrow$ Yahoo $\rightarrow$ DBpedia
-VI-B Implementation details
-Our example encoder is a pretrained BERTBASE model [28].
-The parameter size of our model is 109M.
-For learning rates, $\alpha=1e^{-3}$ and $\beta=3e^{-5}$.
-The training batch size is 16.
-The number of mini-batches in each episode, $m$ = 5.
-The label budget $B_{A}$ = 2000 examples per task and the memory budget $B_{\mathcal{M}}$ is 5 per class, i.e., $n_{s}=5$.
-For textual augmentation, we swap words randomly as the weak augmentation.
-And, we apply the combination of swapping words randomly, deleting words randomly and substituting words by WordNet’s synonym, as the strong augmentation.
-We use nplug [^1] https://github.com/makcedward/nlpaug 1.1.10, a Python package to implement augmentations.
-All models are executed on Linux platform with 8 Nvidia Tesla A100 GPU and 40 GB of RAM.
-All experiments are performed using PyTorch [^2] https://pytorch.org/ 1.10.0+cu113 [29].
-VI-C Baselines
-Baseline model: (a) MAML-SEQ: online FOMAML algorithm.
-(b) OML-ER [8]: OML with 5% episodic experience replay rate [^3] The default 1% rate shows bad performance.
-(c) C-MAML [^4] Being modified for NLP tasks [4]: OML with Meta & CL objective alignment + Reservoir Sampling.
-(d) Meta-CAL (ours.): OML with Meta & CL objective alignment + consistency regularization + reservoir sampling.
-(e) FULL: supervised C-MAML, trained on full datasets.
-Memory sample selection: (f) Prototype [11]: select representative samples that are closest to dynamically updated prototypes in representation space.
-(g) Ring Buffer [30]: use a ‘First-In, First-Out’ scheme to update buffer.
-(h) Reservoir Sampling [3]: randomly select data with an equal selection probability.
-VI-D Evaluation Metrics
-Based on the prior work [31], we use three evaluation metrics, i.e., accuracy, backward transfer and forward transfer.
-Let $R_{k,i}$ be the macro-averaged accuracy evaluated on the test set of the $i$ task after sequentially learning $k$ tasks:
-Accuracy
-$\mathrm{ACC}_{k}=\frac{1}{k}\sum^{k}_{i=1}R_{k,i}$
-Backward transfer
-$\mathrm{BWT}_{k}=\frac{1}{k-1}\sum^{k-1}_{i=1}(R_{k,i}-R_{i,i})$
-Forward transfer
-$\mathrm{FWT}_{k}=\frac{1}{k-1}\sum^{k}_{i=2}(R_{i,i}-R_{0,i}$
-TABLE III: Comparison of different AL strategies. We report the average accuracy of training sequence Yelp $\rightarrow$ AGNews $\rightarrow$ DBpedia $\rightarrow$ Amazon $\rightarrow$ Yahoo.
+FULL [^3] Upper bound performance 64.1 $\pm$ 1.9
+The result shows that our method yields comparable results to the FULL baseline, which is trained on more than 10,000 labeled samples per task.
+This result indicates our method can effectively select 2,000 informative samples from 10,000 unlabeled data to maximize model performance.
+Compared to other baselines with the same MAML framework, our method obtains the highest average accuracy among the four commonly-used AL strategies, demonstrating its robustness to different AL approaches.
+It also indicates Meta-CAL has a great ability of preventing catastrophic forgetting after sequentially learning five tasks.
+We also perform paired t-tests.
+Our model with default setting is significantly better than C-MAML with p-values $<$ 0.04 for all four AL strategies.
+We also compare AL strategies for the proposed model, Random is significantly better than other strategies with p-values $<$ 0.03.
+As for the memory sample selection schemes, Ring Buffer outperforms Reservoir Sampling by less than 1% in RAND and REP.
+Whereas, Reservoir Sampling exhibits smaller standard deviations, indicating its robustness to training set orders.
+Prototype Sampling has the worst performance.
+We perform further analysis in §7.2.
+Table 3: Comparison of Different AL Strategies
 AL Accuracy AVG.
@@
 66.5
+44footnotetext: We report the average accuracy of training sequence Yelp $\rightarrow$ AGNews $\rightarrow$ DBpedia $\rightarrow$ Amazon $\rightarrow$ Yahoo.
+Table 2 and Table 3 demonstrate that Random is the best AL strategy for our model.
+As shown in Table 3, for both Diversity- and Representative-based AL strategies, KMeans is the most effective approach.
+Uncertainty-based methods comparatively perform poorly.
+CL models aim to revisit past samples to enhance generalization.
+However, replaying and annotating uncertain samples may not sufficiently improve generalization capabilities, which can hinder effective knowledge consolidation and consequently harms model performance.
+7 Further Analysis
+We use the training set order [^5] The performance on this order is close to the average performance.
+Yelp $\rightarrow$ AGNews $\rightarrow$ DBpedia $\rightarrow$ Amazon $\rightarrow$ Yahoo to perform further analysis.
+In this section, we denote RAND, REP, DIV and UNC as Random, Representative (KMeans), Diversity (KMeans) and Uncertainty (Least-confidence), respectively.
+7.1 Stability & Plasticity
+We test the performance of different AL strategies in terms of stability and plasticity.
+(a) RAND
+(b) REP (KMeans)
+(c) DIV (KMeans)
+(d) UNC (Least-confidence)
+Figure 1: Per task accuracy at different learning stages.
+The dark color indicates high accuracy.
+From left to right, the color for each task progressively fades away, indicating forgetting happens while learning more tasks.
+Note that Yelp and Amazon are from the same domain (sentiment analysis).UNC shows a lighter color compared to other AL methods, indicating a higher degree of forgetting.
+The red-outlined box shows the accuracy on AGNews (Task 2) after learning DBpedia (Task 3).
+Figure 2: BWT and FWT for different AL strategies at each learning stage.
+Stability: We employ the BWT metric as an indicator of catastrophic forgetting, evaluating the impact of the updated model parameters on the performance across all previously learned tasks.
+The negative BWT values indicate forgetting.
+As shown in Fig.
+2, RAND shows the least forgetting, which indicates the best stability on past tasks.
+However, all methods exhibit a large forgetting after sequentially learning Task 3.
+Task 3 (DBpedia, 14 classes) has a relatively large label space compared to Task 1 (Yelp, 5 classes) and Task 2 (AGNews, 4 classes).
+As shown in Fig.
+1, the accuracy on DBpedia exceeds the accuracy on other tasks.
+Since MAML learns high-quality reusable features for fast adaptation [30], it tends to find shared representations that specifically benefit tasks with large label spaces.
+Plasticity: A positive FWT score indicates the model’s ability to leverage knowledge from previously learned tasks, facilitating zero-shot learning and efficient adaptation to the new task.
+As shown in Fig.
+2, RAND, REP, and DIV show positive FWT values after training Task 4.
+This observation indicates that successful forward knowledge transfer from the previously learned tasks occurs.
+Fig.
+1 further supports that this transfer mainly occurs within the same domain, demonstrating effective domain-incremental learning capabilities.
+Specifically, the proposed method leverages prior knowledge acquired from a familiar domain to facilitate efficient adaptation to new tasks within that domain.
+Overall, RAND shows the best stability while REP shows the best performance in plasticity.
+7.2 Memory Insight
+The level of generalization can be inferred from data dispersion. In this section, we investigate the effect of memory samples resulting from active learning and memory sample selection strategies.
+(a) After Learning Task 1
+(b) After Learning Task 2
+(c) After Learning Task 3
+(d) After Learning Task 4
+(e) After Learning Task 5
+Figure 3: T-SNE visualization of memory samples at different learning stages using training set order Yelp $\rightarrow$ AGNews $\rightarrow$ DBpedia $\rightarrow$ Amazon $\rightarrow$ Yahoo.
+The black-circled data points belong to the last task.
+Data points with darker colors represent samples from earlier tasks, except (d) after learning Task 4.
+Task 4 has the same domain as Task 1.
+Hence, data points with darker colors are belong to the latest task in (d).
+Different learning stages: Fig.
+3 presents the T-SNE visualization of memory samples at different learning stages.
+As show in Figure 3a to 3b, when the proposed model learns a small number of tasks or encounters small label spaces, it focuses on ensuring intra-task generalization.
+However, as the number of seen tasks increases, the model shifts its focus towards ensuring inter-task generalization as shown in 3d to 3e.
+Consequently, we observe a phenomenon wherein the memory samples from the last tasks cluster together.
 (a) Reservoir-RAND(65.8%)
@@
 The clustering of memory samples from the last task suggests model focuses more on inter-task generalization rather than on intra-task generalization.
-VI-E Results
-In Table I, we compare various models with four AL strategies.
-These strategies include Random (denoted as RAND), Representative via KMeans (denoted as REP), Diversity via KMeans (denoted as DIV), and Uncertainty via Least-confidence (denoted as UNC).
-The label budget $B_{A}$ = 2000 examples per task.
-Each record is the average of three best results from five runs.
-Save ratio is the ratio of saved samples to the total number of seen samples.
-The result shows that our method yields comparable results to the FULL baseline, which is trained on more than 10,000 labeled samples per task.
-This result indicates our method can effectively select 2,000 informative samples from 10,000 unlabeled data to maximize model performance.
-Compared to other baselines with the same MAML framework, our method obtains the highest average accuracy among the four commonly-used AL strategies, demonstrating its robustness to different AL approaches.
-It also indicates Meta-CAL has a great ability of preventing catastrophic forgetting after sequentially learning five tasks.
-We also perform paired t-tests.
-Our model with default setting is significantly better than C-MAML with p-values $<$ 0.04 for all four AL strategies.
-We also compare AL strategies for the proposed model, Random is significantly better than other strategies with p-values $<$ 0.03.
-As for the memory sample selection schemes, Ring Buffer outperforms Reservoir Sampling by less than 1% in RAND and REP.
-Whereas, Reservoir Sampling exhibits smaller standard deviations, indicating its robustness to training set orders.
-Prototype Sampling has the worst performance.
-We perform further analysis in §VII-A.
-Both Table I and Table III demonstrate that Random is the best AL strategy for our model.
-As shown in Table III, for both Diversity- and Representative-based AL strategies, KMeans is the most effective approach.
-The uncertainty-based methods comparatively perform poorly.
-CL models intend to revisit past samples that can bring generalization.
-However, annotating and replaying uncertain samples cannot bring enough generalization, which is not beneficial for knowledge consolidation.
-Hence, it harms model performance.
-VII Further Analysis
-We use the training set order [^5] The performance on this order is close to the average performance.
-Yelp $\rightarrow$ AGNews $\rightarrow$ DBpedia $\rightarrow$ Amazon $\rightarrow$ Yahoo to perform further analysis.
-In this section, we denote RAND, REP, DIV and UNC as Random, Representative (KMeans), Diversity (KMeans) and Uncertainty (Least-confidence), respectively.
-(a) After Learning Task 1
-(b) After Learning Task 2
-(c) After Learning Task 3
-(d) After Learning Task 4
-(e) After Learning Task 5
-Figure 2: T-SNE visualization of memory samples at different learning stages using training set order Yelp $\rightarrow$ AGNews $\rightarrow$ DBpedia $\rightarrow$ Amazon $\rightarrow$ Yahoo.
-The black-circled data points belong to the last task.
-Data points with darker colors represent samples from earlier tasks, except (d) after learning Task 4.
-Task 4 has the same domain as Task 1.
-Hence, data points with darker colors are belong to the latest task in (d).
-(a) RAND
-(b) REP (KMeans)
-(c) DIV (KMeans)
-(d) UNC (Least-confidence)
-Figure 3: Per task accuracy at different learning stages.
-The dark color indicates high accuracy.
-From left to right, the color for each task progressively fades away, indicating forgetting happens while learning more tasks.
-Note that Yelp and Amazon are from the same domain (sentiment analysis).UNC shows a lighter color compared to other AL methods, indicating a higher degree of forgetting.
-The red-outlined box shows the accuracy on AGNews (Task 2) after learning DBpedia (Task 3).
-VII-A Memory Insight
-The level of generalization can be inferred from data dispersion. In this section, we investigate the effect of memory samples resulting from active learning and memory sample selection strategies.
-AL strategies
-As shown in Fig.
-1a to 1d, memory data in RAND show a good data dispersion within the last task (i.e., intra-task generalization) and across multiple tasks (i.e., inter-task generalization), resulting in the best accuracy.
+AL strategies: As shown in Fig.
+4a to 4d, memory data in RAND show a good data dispersion within the last task (i.e., intra-task generalization) and across multiple tasks (i.e., inter-task generalization), resulting in the best accuracy.
 In contrast, UNC in Fig.
-1d shows subpar performance in inter-task generalization.
+4d shows subpar performance in inter-task generalization.
 The result validates UNC’s incapability to achieve generalization and preserve knowledge.
@@
 Therefore, it is important to find an optimal balance between representativeness and diversity to ensure generalization.
-Memory sample selection methods
-In Fig.
-1a, 1e and 1f, the choice of memory sample selection method also affects the model’s performance.
+Memory sample selection methods: In Fig.
+4a, 4e and 4f, the choice of memory sample selection method also affects the model’s performance.
 Prototype Sampling selects representative memory samples, potentially resulting in relatively less sparsity across multiple tasks.
@@
 This validates the superiority of our method in addressing confusion between old and new tasks.
-Different learning stages
-Fig.
-2 presents the T-SNE visualization of memory samples at different learning stages.
-As show in Figure 2a to 2b, when the proposed model learns a small number of tasks or encounters small label spaces, it focuses on ensuring intra-task generalization.
-However, as the number of seen tasks increases, the model shifts its focus towards ensuring inter-task generalization as shown in 2d to 2e.
-Consequently, we observe a phenomenon wherein the memory samples from the last tasks cluster together.
-VII-B Stability & Plasticity
-Next we test the performance of different AL strategies in terms of stability and plasiticity.
-(a) BWT (Stability)
-(b) FWT (Plasticity)
-Figure 4: BWT and FWT for different AL strategies at each learning stage.
-Stability
-The negative BWT values indicate forgetting.
-As shown in Fig.
-4a, RAND shows the least forgetting, which indicates the best stability on past tasks.
-However, all methods exhibit a large forgetting after sequentially learning Task 3.
-Task 3 (DBpedia, 14 classes) has a relatively large label space compared to Task 1 (Yelp, 5 classes) and Task 2 (AGNews, 4 classes).
-As shown in Fig.
-3, the accuracy on DBpedia exceeds the accuracy on other tasks.
-Since MAML learns high-quality reusable features for fast adaptation [32], it tends to find shared representations that specifically benefit tasks with large label spaces.
-Plasticity
-A positive FWT demonstrates the model’s ability of zero-shot learning.
-Task 4 (Amazon) are from the same domain as Task 1 (Yelp).
-RAND, REP and DIV have a positive FWT value after learning Amazon.
-The result suggests a positive forward knowledge transfer to the task from the same domain.
-Moreover, it demonstrates the effectiveness of our model in domain-incremental learning when employed with these three AL strategies.
-Overall, RAND shows the best stability while REP shows the best performance in plasticity.
-VII-C Effect of Augmentations
+7.3 Effect of Augmentations
 We conduct an ablation study to analyze textual augmentations and their key roles in Meta-CAL.
-We examine the effect of inner-loop augmentation and outer loop augmentation in Table IV.
-The query samples in the outer-loop constrain the model behaviour.
-We hypothesize that learning sufficient generalizability from query samples helps retain knowledge and prevent catastrophic forgetting.
-The results validate this hypothesis, as we observe a significant gain by using outer-loop augmentation to enhance generalization.
-It improves accuracy by approximately 10% when data availability and label budget are extremely limited.
-Inner-loop augmentation works well in the extreme case.
-Nevertheless, the combination of inner- and outer-loop augmentations still improves accuracy.
-TABLE IV: Ablation study with different textual augmentation modules of Meta-CAL
+We examine the effect of inner-loop augmentation and outer loop augmentation in Table 4.
+The meta samples in the outer-loop constrain the model behaviour.
+Table 4: Ablation Study with Different Augmentation Modules
 AL strategy: 5-shot 1-shot
@@
 $\pm$ 0.9
-VII-D Annotation Budgets
+We hypothesize that if the meta-samples exhibit sufficient generalization capabilities, they can facilitate effective knowledge retention from prior tasks, consequently mitigating the issue of catastrophic forgetting.
+It is noteworthy that the meta-samples acquired through the combined effect of active learning acquisition and memory sample selection.
+We further enhance their generalization capabilities through data augmentation.
+The results validate this hypothesis, as we observe a significant gain by using outer-loop augmentation to enhance generalization.
+It improves accuracy by approximately 10% when data availability and label budget are extremely limited.
+While inner-loop augmentation proves effective in extreme cases, it might not be as advantageous as outer-loop augmentation in non-extreme scenarios.
+Nevertheless, the combination of inner- and outer-loop augmentations still significantly improves accuracy.
+7.4 Annotation Budgets
 As shown in Fig.
@@
 Figure 5: Performance on different annotation budgets.
-VII-E Memory Budgets
+7.5 Memory Budgets
 We evaluate memory efficiency of our model.
-Table V compares the performance of three models with the best average accuracy, i.e., our model with Reservoir sampling, our model with Ring Buffer and C-MAML.
-While all these models use the MAML framework, only the Meta-CAL models employ data augmentation to enhance generalization.
+Table 5 compares the performance of three models with the best average accuracy, i.e., our model with Reservoir sampling, our model with Ring Buffer and C-MAML.
+While all these models use the MAML framework, only the Meta-CAL models employ consistency regularization to enhance generalization.
 The result shows Meta-CAL models outperform C-MAML in all three cases.
@@
 Reservoir can attain more than 50% accuracy by saving only one sample per seen class.
-As a result, it indicates that data augmentation improves performance.
-TABLE V: Accuracy on different memory budgets. The AL strategy is RAND.
+As a result, it also indicates that consistency regularization through data augmentation improves performance.
+Table 5: Accuracy on different memory budgets. The AL strategy is RAND.
 Memory Budget Meta-CAL C-MAML
@@
 $\pm$ 0.6
-Specifically, our model dynamically selects worthwhile unlabeled data to annotate, thereby preventing catastrophic forgetting and maximizing model performance.
-VIII Conclusion
-This paper considers a realistic continual learning scenario with limited annotation budgets, namely continual active learning.
-We devise a novel method to address this problem.
-Our method dynamically selects worthwhile unlabeled data for annotation to maximize model performance and address continual learning problems.
-Specifically, it applies MAML with memory replay and consistency regularization to enable efficient learning with limited memory and annotation budgets while mitigating catastrophic forgetting.
-In experiments, we analyzed the effect of various active learning strategies and memory sample selection methods on model performance.
-The experimental results shows our method achieves a great performance in a 5-shot continual active learning setup.
-Furthermore, we demonstrate that random sampling is still the best default strategy for both active learning and memory sample selection to solve 5-shot continual active learning problems.
-However, we only evaluate our method in a 5-shot continual active learning scenario.
-In the future work, we can extend our evaluation to a more realistic scenario where the amount of labeled data is {100, 1000, 5000}.
-In addition, the annotation budget is allocated equally for each task.
+8 Conclusion
+This paper considers a realistic continual learning scenario, namely few-shot continual active learning.
+To address this resource-constrained continual learning problems, we propose a novel method that employing meta-learning and active learning techniques, called Meta-Continual Active Learning.
+Specifically, our model dynamically queries worthwhile unlabeled data for annotation and reformulate meta-objective with experience rehearsal and consistency regularization.
+In such a way, the proposed method is able to prevent catastrophic forgetting and improve generalization capability in a low-resource scenario.
+We conduct extensive experiments on benchmark text classification datasets in a 5-shot continual active learning setting.
+The results show the robustness of the proposed method.
+However, we only evaluate our method in a 5-shot case.
+In the future work, we can extend our evaluation to a more realistic scenario where the amount of labeled data is {100, 1000, 5000} and extend to other NLP tasks, e.g., language model training, text generation, and knowledge base enrichment.
+Furthermore, the annotation budget is allocated equally for each task.
 The annotation budget allocation strategies can be further studied.
+Appendix A Supplemental Material
+Table 6: Accuracy of Four Training Set Orders and Standard Deviation
+AL CL Method Order 1 Order 2 Order 3 Order 4 Average
+RAND MAML-SEQ $4.1\pm 0.3$ $10.1\pm 0.2$ $14.8\pm 3.6$ $4.4\pm 3.9$ $8.4\pm 5.1$
+OML-ER $60.2\pm 3.0$ $59.7\pm 2.1$ $56.4\pm 2.7$ $60.4\pm 4.0$ $59.2\pm 1.9$
+C-MAML $61.2\pm 2.1$ $65.1\pm 2.8$ $65.8\pm 1.0$ $59.8\pm 2.3$ $63.0\pm 2.9$
+Ours.(PROT) $54.7\pm 1.8$ $61.8\pm 1.6$ $62.1\pm 1.8$ $52.4\pm 2.7$ $57.6\pm 4.9$
+Ours.(RB) $66.4\pm 0.5$ $68.1\pm 0.6$ $69.2\pm 0.2$ $63.8\pm 0.2$ $66.9\pm 2.4$
+Ours.(RESV) $65.7\pm 0.3$ $68.1\pm 1.2$ $68.1\pm 1.2$ $64.2\pm 0.6$ $66.5\pm 1.9$
+MAML-SEQ $10.3\pm 7.1$ $12.2\pm 1.0$ $6.6\pm 2.7$ $1.4\pm 0.1$ $7.63\pm 4.8$
+OML-ER $55.5\pm 3.1$ $56.4\pm 1.2$ $58.2\pm 3.4$ $59.3\pm 3.2$ $57.4\pm 1.7$
+REP C-MAML $58.3\pm 0.7$ $62.0\pm 2.6$ $64.4\pm 1.8$ $58.9\pm 3.1$ $60.9\pm 2.8$
+(KMeans) Ours.(PROT) $58.0\pm 2.8$ $62.3\pm 1.3$ $59.2\pm 3.3$ $51.8\pm 0.6$ $57.8\pm 4.4$
+Ours.(RB) $65.3\pm 1.0$ $67.3\pm 0.6$ $68.7\pm 1.4$ $63.7\pm 1.5$ $66.3\pm 2.2$
+Ours.(RESV) $65.2\pm 0.5$ $67.5\pm 0.9$ $67.3\pm 0.2$ $63.5\pm 1.2$ $65.9\pm 1.9$
+MAML-SEQ $7.3\pm 2.2$ $9.6\pm 1.6$ $10.5\pm 6.3$ $1.5\pm 0.1$ $7.23\pm 4.1$
+OML-ER $54.8\pm 1.8$ $54.5\pm 1.2$ $59.1\pm 2.8$ $60.8\pm 3.1$ $57.3\pm 3.1$
+DIV C-MAML $60.9\pm 1.7$ $59.1\pm 2.1$ $60.6\pm 1.7$ $59.1\pm 2.1$ $59.9\pm 1.0$
+(KMeans) Ours.(PROT) $55.5\pm 3.2$ $63.4\pm 1.4$ $59.7\pm 1.1$ $52.6\pm 4.6$ $57.8\pm 4.7$
+Ours.(RB) $63.7\pm 0.4$ $66.8\pm 1.5$ $66.8\pm 0.3$ $61.7\pm 2.1$ $64.8\pm 2.5$
+Ours.(RESV) $64.7\pm 1.1$ $65.7\pm 0.6$ $66.1\pm 0.2$ $63.1\pm 2.3$ $64.9\pm 1.3$
+MAML-SEQ $4.2\pm 2.0$ $9.2\pm 1.8$ $9.6\pm 4.3$ $1.6\pm 0.4$ $6.15\pm 3.9$
+OML-ER $47.1\pm 1.5$ $51.9\pm 2.8$ $46.7\pm 7.6$ $44.0\pm 1.5$ $47.4\pm 3.3$
+UNC C-MAML $50.9\pm 3.3$ $49.2\pm 1.5$ $53.1\pm 2.5$ $48.4\pm 0.4$ $50.4\pm 2.1$
+(L-CONF) Ours.(PROT) $37.5\pm 6.5$ $49.5\pm 0.9$ $44.6\pm 6.0$ $36.3\pm 3.5$ $42.0\pm 6.2$
+Ours.(RB) $47.4\pm 3.6$ $53.9\pm 1.1$ $53.7\pm 2.0$ $47.8\pm 1.4$ $50.7\pm 3.6$
+Ours.(RESV) $53.3\pm 3.8$ $54.4\pm 0.5$ $56.3\pm 1.3$ $49.8\pm 0.8$ $53.5\pm 2.7$
+REP(Means) Ours.(RESV) $57.5\pm 0.9$ $58.1\pm 0.4$ $60.6\pm 1.2$ $55.5\pm 1.0$ $57.9\pm 2.1$
+DIV(Means) $50.0\pm 1.5$ $50.5\pm 2.8$ $52.5\pm 0.4$ $49.6\pm 4.7$ $50.7\pm 1.3$
+UNC(MARG) $58.4\pm 1.8$ $58.0\pm 0.9$ $59.1\pm 2.1$ $49.2\pm 1.4$ $56.2\pm 4.7$
+UNC(ENT) $51.2\pm 0.3$ $49.2\pm 1.4$ $54.6\pm 2.3$ $48.5\pm 4.1$ $50.9\pm 2.7$
+FULL C-MAML $64.5\pm 1.4$ $62.2\pm 1.3$ $66.5\pm 1.0$ $63.2\pm 1.7$ $64.1\pm 1.7$